### v4

# Conjugate Direction Methods Under Inconsistent Systems

Abstract
Since the development of the conjugate gradient (CG) method in 1952 by Hestenes and Stiefel, CG, has become an indispensable tool in computational mathematics for solving positive definite linear systems. On the other hand, the conjugate residual (CR) method, closely related CG and introduced by Stiefel in 1955 for the same settings, remains relatively less known outside the numerical linear algebra community. Since their inception, these methods – henceforth collectively referred to as conjugate direction methods – have been extended beyond positive definite to indefinite, albeit consistent, settings. Going one step further, in this paper, we investigate theoretical and empirical properties of these methods under inconsistent systems. Among other things, we show that small modifications to the original algorithms allow for the pseudo-inverse solution. Furthermore, we show that CR is essentially equivalent to the minimum residual method, proposed by Paige and Saunders in 1975, in such contexts. Lastly, we conduct a series of numerical experiments to shed lights on their numerical stability (or lack thereof) and their performance for inconsistent systems. Surprisingly, we will demonstrate that, unlike CR and contrary to popular belief, CG can exhibit significant numerical instability, bordering on catastrophe in some instances.
1 Introduction
Consider the following linear system,
$\mathbf{Ax}=\mathbf{b},$ (1)
where $\mathbf{A}\in\mathbb{R}^{d\times d}$ is a symmetric matrix and $\mathbf{b}\in\mathbb{R}^{d}$. In 1952, Hestenes and Stiefel [22] developed the celebrated conjugate gradient (CG) method to solve 1 when $\mathbf{A}$ is positive definite (PD). Since then, CG has become a cornerstone and essential tool in many areas of applied mathematics, statistics, machine learning, and numerical optimization. For example, in solving for numerical solutions of partial differential equations, the discretization of a differential operator often leads to a positive definite matrix, making CG the go-to method. In statistical modeling and machine learning, CG is frequently the primary solver in methods involving kernel regression. In numerical optimization, the classical Newton-CG method [33] has been the canonical example of truncated Newton-type methods for large-scale problems. A closely related technique to CG is the conjugate residual (CR) method, which was developed by Stiefel in 1955 [43] for the same PD settings [^4] In fact, without explicitly mentioning the name, the structure of the method is developed in the original paper of Hestenes and Stiefel [22].. However, despite its profound connections to CG, the CR method remains relatively less known outside the numerical linear algebra community. For example, to our knowledge, the only attempt at employing CR within Newton-type optimization methods is Dahito et al [6].
Both these methods, henceforth collectively referred to as conjugate direction (CD) methods [^5] Precisely speaking, the conjugate direction approach serves as a framework for deriving various methods, including CG and CR. However, in this paper, when we refer to CD, we specifically mean these two particular instances., are examples of Krylov subspace algorithms for solving 1 when $\mathbf{A}$ is PD. However, advancements in science and engineering have led to a growing need for iterative solvers that can effectively address a broader range of challenges beyond PD systems. This includes scenarios where the matrix $\mathbf{A}$ is symmetric, but singular and/or indefinite, as well as inconsistent settings, namely $\mathbf{b}\not\in\text{Range}(\mathbf{A})$, e.g., [44, 38, 36, 31]. As a result, the popularity of CD, especially CG, to solve problems under these conditions continues to extend. There are studies [12, 29, 30, 8, 9, 10, 15, 42] providing comprehensive treatment of CD beyond positive definite systems. However, all these results operate under the assumption that the system is consistent, i.e. $\mathbf{b}\in\text{Range}(\mathbf{A})$.
In contrast, the minimum residual [^6] We would like to caution the reader that Algorithm 5.3 in the book [40] is called Minimal Residual Iteration which is entirely unrelated to MINRES. (MINRES) method by Paige and Saunders [34] extends naturally to all systems involving symmetric matrices, without relying on a consistency assumption and has shown great success in numerical optimization in recent years, e.g., [36, 24, 27, 28]. In comparison to CG or CR, however, MINRES is a significantly more complex algorithm and entails more intricate implementations. Therefore, understanding the properties and the behavior of the CD methods in inconsistent settings can be of great interest. Before delving any deeper, it is essential to note that our objective in this paper is not to devise novel variants of CG or CR that can extend to inconsistent systems. Instead, our emphasis is on exploring the theoretical and empirical behavior of the original methods in such situations.
Contributions
For symmetric, but inconsistent, systems, the contributions of this paper can be informally summarized as follows:
  1. (i)
    We first show that the classical CG is fundamentally unable to recover any solution of the normal equation (Theorem 1). Subsequently, we provide a modification (Algorithm 3) that ensures convergence to the pseudo-inverse solution (Theorem 2).
  2. (ii)
    We provide a variant of CR (Algorithm 4) that converges to the pseudo-inverse solution (Theorem 4). We also establish that, in essence, CR and MINRES are equivalent (Theorem 5).
  3. (iii)
    We numerically demonstrate that, beyond typical finite precision errors, CG–unlike CR–can exhibit borderline catastrophic instability in inconsistent settings. These instabilities can persist even for PD matrices with a few small eigenvalues.
Table 1 highlights some of the similarities and differences between vanilla CG and CR at termination.
Consistent: $\mathbf{b}\in\text{Range}(\mathbf{A})$ Inconsistent: $\mathbf{b}\notin\text{Range}(\mathbf{A})$
CG CR CG CR
$\mathbf{x}=\mathbf{x}^{\star}$
(Theorem 3)
$\mathbf{x}=\mathbf{x}^{+}$
(Fact 1) (w. Algorithm 3) (w. Algorithm 4)
$\mathbf{r}=\mathbf{0}$
$\mathbf{Ar}=\mathbf{0}$
(Theorem 1) (Theorem 3)
$\mathbf{p}=\mathbf{0}$
(Lemma 1) (Lemma 3)
$\mathbf{Ap}=\mathbf{0}$ (Fact 2)
(Lemma 1) (Lemma 3)
Table 1: Similarities and differences between CG (Algorithm 1) and CR (Algorithm 2) at termination. Here, $\mathbf{x}$ is the final iterate, $\mathbf{r}=\mathbf{b}-\mathbf{Ax}$ is the residual at $\mathbf{x}$, $\mathbf{p}$ is the update direction within the CD methods, $\mathbf{x}^{+}$ and $\mathbf{x}^{\star}$ denote the pseudo-inverse solution and a normal solution with $\mathbf{y}\neq\mathbf{0}$ from Definition 1, respectively. All quantities are at the termination of the respective algorithm.
The rest of this paper is organized as follows. We end this section by introducing the notation and definitions that are used throughout this work. We first review the CD methods for consistent systems in Section 2. Subsequently, we shift our focus to the central theme of our paper and present theoretical results for inconsistent systems in Section 3.2.2. Finally, in Section 4, we provide numerical experiments to verify our theoretical results as well as to shed light on the numerical performance of CD methods in various situations.
Notations and Definitions
We denote scalars, vectors, and matrices as lower-case, bold lower-case, and bold upper-case letters, e.g., $\alpha$, $\mathbf{b}$ and $\mathbf{A}$, respectively. For two vectors $\mathbf{a}\in\mathbb{R}^{d}$ and $\mathbf{b}\in\mathbb{R}^{d}$, their Euclidean inner product is denoted as $\left\langle\mathbf{a},\mathbf{b}\right\rangle=\mathbf{a}^{\top}\mathbf{b}$. Similarly, for a symmetric matrix $\mathbf{A}\in\mathbb{R}^{d\times d}$, the product $\mathbf{a}^{\top}\mathbf{Ab}$ (not necessarily an inner-product) is denoted by $\left\langle\mathbf{a},\mathbf{b}\right\rangle_{\mathbf{A}}$. The norm $\left\|\mathbf{a}\right\|=\sqrt{\left\langle\mathbf{a},\mathbf{a}\right\rangle}$ is the Euclidean norm, and for a PD matrix $\mathbf{A}$, we define the energy norm $\|\mathbf{a}\|_{\mathbf{A}}=\sqrt{\left\langle\mathbf{a},\mathbf{A}\mathbf{a}\right\rangle}$. We write $\mathbf{A}\succeq\mathbf{0}$ and $\mathbf{A}\succ\mathbf{0}$ to indicate $\mathbf{A}$ is positive semi-definite (PSD) and PD, respectively. The Moore-Penrose pseudo-inverse of $\mathbf{A}$ is denoted by $\mathbf{A}^{\dagger}$.
For a symmetric matrix $\mathbf{A}$ and $\mathbf{b}\not\in\text{Range}(\mathbf{A})$, we denote the orthogonal decomposition of $\mathbf{b}$ with respect to $\mathbf{A}$ as
$\mathbf{b}\triangleq\tilde{\mathbf{b}}+\mathbf{b}^{\perp},\quad\text{where}\quad\tilde{\mathbf{b}}\in\text{Range}(\mathbf{A}),\quad\text{and}\quad\mathbf{b}^{\perp}\in\text{Null}(\mathbf{A}).$ (2)
So, for example, for consistent systems, we have $\mathbf{b}^{\perp}=\mathbf{0}$. By CG$(\mathbf{A},\mathbf{b})$, CR$(\mathbf{A},\mathbf{b})$ and MINRES$(\mathbf{A},\mathbf{b})$ we mean using CG, CR and MINRES to solve for a least-squares problem or a linear system with $\mathbf{A}$ and $\mathbf{b}$, respectively. Furthermore, under inconsistent systems, we will often compare between two settings, say, CG$(\mathbf{A},\mathbf{b})$ and its corresponding consistent case CG$(\mathbf{A},\tilde{\mathbf{b}})$, where $\tilde{\mathbf{b}}$ is as in the orthogonal decomposition of $\mathbf{b}$ above. We denote $\alpha_{k-1}$, $\mathbf{r}_{k-1}$, etc, to be the terms generated for solving the former system and use $\tilde{\alpha}_{k-1}$, $\tilde{\mathbf{r}}_{k-1}$ etc, to denote the terms generated for solving the latter system. The residual vector at iteration $k$ is denoted by $\mathbf{r}_{k}=\mathbf{b}-\mathbf{Ax}_{k}$.
Definition 1 (Normal Solutions and the Pseudo-inverse Solution).
Any solution to the normal equation $\mathbf{A}^{\top}\mathbf{A}\mathbf{x}=\mathbf{A}^{\top}\mathbf{b}$, namely $\mathbf{x}^{\star}\triangleq\mathbf{A}^{\dagger}\mathbf{b}+(\mathbf{I}-\mathbf{A}^{\dagger}\mathbf{A})\mathbf{y}$ for some $\mathbf{y}\in\mathbb{R}^{d}$, is referred to as a normal solution. In particular, the pseudo-inverse solution is given by $\mathbf{x}^{+}\triangleq\mathbf{A}^{\dagger}\mathbf{b}$, i.e., when $\mathbf{y}=\mathbf{0}$.
The Krylov subspace of degree $k\geq 1$, generated by $\mathbf{A}$ and $\mathbf{v}$, is denoted as
$\mathcal{K}_{k}(\mathbf{A},\mathbf{v})=\text{Span}\{\mathbf{v},\mathbf{Av},\cdots,\mathbf{A}^{k-2}\mathbf{v},\mathbf{A}^{k-1}\mathbf{v}\}.$
Recall that the grade of $\mathbf{v}$ with respect to $\mathbf{A}$ is defined as follows [1].
Definition 2 (Grade of $\mathbf{v}$ with respect to $\mathbf{A}$).
The grade of $\mathbf{v}$ with respect to $\mathbf{A}$ is the positive integer $g\in\mathbb{N}$ such that
$\dim(\mathcal{K}_{k}(\mathbf{A},\mathbf{v}))=\begin{cases}k,&k\leq g,\\ g,&k>g.\end{cases}$
Remark 1 (Initialization).
In this paper, we assume $\mathbf{x}_{0}=\mathbf{0}$. Thus, the initial residual and the Krylov subspace become $\mathbf{r}_{0}=\mathbf{b}-\mathbf{Ax}_{0}=\mathbf{b}$ and $\mathcal{K}_{k}(\mathbf{A},\mathbf{b})=\text{Span}(\mathbf{b},\cdots,\mathbf{A}^{k-1}\mathbf{b})$ respectively. Since extensions to arbitrary initialization are straightforward, we exclude them for the sake of brevity.
Remark 2 (Symmetric Matrix and Non-zero Right-hand side).
Throughout the paper, for simplicity, we consider systems involving real symmetric matrices. We note that it is straightforward to adjust all statements for Hermitian matrices. Furthermore, we always assume $\mathbf{b}\neq\mathbf{0}$ to avoid triviality.
For indefinite matrices, we may encounter a non-zero vector $\mathbf{x}\notin\text{Null}(\mathbf{A})$ for which $\left\langle\mathbf{x},\mathbf{Ax}\right\rangle=0$, i.e., $\mathbf{x}\perp\mathbf{Ax}$. Such a vector is henceforth referred to as a zero-curvature direction.
Definition 3 (Zero-curvature Direction).
For a given indefinite matrix $\mathbf{A}$, a non-zero vector $\mathbf{x}\notin\text{Null}(\mathbf{A})$ is a zero-curvature direction if $\left\langle\mathbf{x},\mathbf{Ax}\right\rangle=0$.
2 Consistent Systems: Review
Conjugate direction methods have been extensively studied in the literature for consistent settings where $\mathbf{b}\in\text{Range}(\mathbf{A})$. In this section, we offer a concise overview of these methods in this context and present some results that, while potentially well-known, may not be readily available in a unified source in the literature.
[…]
$\mathbf{x}_{k}=\underset{\mathbf{x}\in\mathcal{K}_{k}(\mathbf{A},\mathbf{b})}{\text{argmin}}\;\frac{1}{2}\left\langle\mathbf{x},\mathbf{A}\mathbf{x}\right\rangle-\left\langle\mathbf{x},\mathbf{b}\right\rangle.$ (3)
Equivalently, one can derive $\mathbf{x}_{k}$ by enforcing the Petrov-Galerkin condition [1], as the vector in the search space $\mathcal{K}_{k}(\mathbf{A},\mathbf{b})$ whose residual $\mathbf{r}_{k}$ is orthogonal to a constraint space, which in the case of CG is taken to be the same as $\mathcal{K}_{k}(\mathbf{A},\mathbf{b})$. When $k=g$, i.e., the grade of $\mathbf{b}$ with respect to $\mathbf{A}$, the solution that minimizes 3 is also the solution to 1.
[…]
$\mathbf{x}_{k}=\underset{\mathbf{x}\in\mathcal{K}_{k}(\mathbf{A},\mathbf{b})}{\text{argmin}}\;\frac{1}{2}\left\langle\mathbf{x},\mathbf{A}\mathbf{x}\right\rangle_{\mathbf{A}}-\left\langle\mathbf{x},\mathbf{b}\right\rangle_{\mathbf{A}}.$ (4)
Consequently, the Petrov-Galerkin condition here involves the same search space as that of CG, namely $\mathcal{K}_{k}(\mathbf{A},\mathbf{b})$, but the constraint space is now $\mathbf{A}\mathcal{K}_{k}(\mathbf{A},\mathbf{b})$. Furthermore, it is easy to see that 4 can be formulated as
$\mathbf{x}_{k}=\underset{\mathbf{x}\in\mathcal{K}_{k}(\mathbf{A},\mathbf{b})}{\text{argmin}}\;\frac{1}{2}\|\mathbf{b}-\mathbf{Ax}\|^{2},$ (5)
highlighting the well-known equivalence between CR and MINRES for when $\mathbf{A}\succ\mathbf{0}$ [14]. We remind that while both CG and CR methods involve one matrix-vector product in each iteration, CR requires one more vector of storage and one more vector update compared with CG. Indeed, in Algorithm 2, after computing $\mathbf{Ar}_{k}$, one can calculate $\mathbf{Ap}_{k}$ as $\mathbf{Ap}_{k}=\mathbf{Ar}_{k}+\beta_{k}\mathbf{Ap}_{k-1}$; see [1] for more details.
Algorithm 1 Conjugate Gradient
1:$\mathbf{A}$ and $\mathbf{b}$
2:$k=1$, $\mathbf{x}_{0}=\mathbf{0}$, $\mathbf{r}_{0}=\mathbf{p}_{0}=\mathbf{b}$
3:while Not Terminated do
4:     $\alpha_{k-1}=\|\mathbf{r}_{k-1}\|^{2}/\left\langle\mathbf{p}_{k-1},\mathbf{Ap}_{k-1}\right\rangle$
5:     $\mathbf{x}_{k}=\mathbf{x}_{k-1}+\alpha_{k-1}\mathbf{p}_{k-1}$
6:     $\mathbf{r}_{k}=\mathbf{r}_{k-1}-\alpha_{k-1}\mathbf{Ap}_{k-1}$
7:     $\beta_{k-1}=\|\mathbf{r}_{k}\|^{2}/\|\mathbf{r}_{k-1}\|^{2}$
8:     $\mathbf{p}_{k}=\mathbf{r}_{k}+\beta_{k-1}\mathbf{p}_{k-1}$
9:     $k=k+1$
10:end while
11:return $\mathbf{x}_{k}$
Algorithm 2 Conjugate Residual
1:$\mathbf{A}$ and $\mathbf{b}$
2:Let $k=1$, $\mathbf{x}_{0}=\mathbf{0}$, $\mathbf{r}_{0}=\mathbf{p}_{0}=\mathbf{b}$
3:while Not Terminated do
4:     $\alpha_{k-1}=\|\mathbf{r}_{k-1}\|^{2}_{\mathbf{A}}/\left\langle\mathbf{p}_{k-1},\mathbf{Ap}_{k-1}\right\rangle_{\mathbf{A}}$
5:     $\mathbf{x}_{k}=\mathbf{x}_{k-1}+\alpha_{k-1}\mathbf{p}_{k-1}$
6:     $\mathbf{r}_{k}=\mathbf{r}_{k-1}-\alpha_{k-1}\mathbf{Ap}_{k-1}$
7:     $\beta_{k-1}=\|\mathbf{r}_{k}\|^{2}_{\mathbf{A}}/\|\mathbf{r}_{k-1}\|^{2}_{\mathbf{A}}$
8:     $\mathbf{p}_{k}=\mathbf{r}_{k}+\beta_{k-1}\mathbf{p}_{k-1}$
9:     $k=k+1$
10:end while
11:return $\mathbf{x}_{k}$
Going beyond PD settings, there have been several efforts in extending the CD methods. However, in all these attempts, the underlying assumption is that the linear system is consistent. The extension to consistent systems involving positive semi-definite (PSD) matrices is relatively straightforward and has been explored in the literature [15, 19, 18]. In the work of Kaasschieter [23], the behavior of preconditioned CG under positive semi-definite, but consistent, systems is studied. The main challenge in extending the CD methods to indefinite settings is the risk of encountering division by zero. In CG, for example, this occurs when $\left\langle\mathbf{p}_{k-1},\mathbf{Ap}_{k-1}\right\rangle=0$ at which point $\alpha_{k-1}$ is no longer well-defined and CG breaks down. This can happen when either $\mathbf{Ap}_{k-1}=\mathbf{0}$ or $\mathbf{Ap}_{k-1}\perp\mathbf{p}_{k-1}$. We will show later that the former case constitutes a breakdown, as it signals the final iteration and CG can be safely terminated (Lemma 1). The latter case, however, is considered an “unlucky breakdown”. For CG applied to indefinite, but consistent, settings, this situation is handled by employing the concept of a hyperbolic pair [12, 29]. Likewise, there are works [8, 9, 10], which developed planar CG, offering an alternative approach for handling unlucky breakdowns within CG for indefinite, but consistent, settings. In indefinite settings, similar breakdowns can occur in CR when $\left\langle\mathbf{r}_{k-1},\mathbf{Ar}_{k-1}\right\rangle=0$ at which point $\beta_{k-1}$ is not well-defined and CR breaks down. This can happen when either $\mathbf{Ar}_{k-1}=\mathbf{0}$ or $\mathbf{Ar}_{k-1}\perp\mathbf{r}_{k-1}$. Again, we will later show that the former case is a breakdown that is theoretically guaranteed by Theorem 3, while the latter scenario is again characterized as an unlucky breakdown.
As mentioned earlier, our goal in this paper is not to design new variants of CG or CR that can extend to indefinite systems. Instead, our emphasis is on exploring the convergent behavior of the original methods in such situations. As a result, to exclude cases where the algorithm encounters an unlucky breakdown in indefinite settings, we make an umbrella assumption that a zero-curvature, as in Definition 3, is not encountered in the algorithm prior to termination.
Assumption 1 (No Unlucky Breakdown).
If $\mathbf{A}$ is indefinite, the CD methods do not encounter a zero-curvature direction as in Definition 3.
Without Assumption 1, the convergence behavior of CG and CR for inconsistent systems can generally be studied based on whether or not an unlucky breakdown occurs during the iterations. In the case of an unlucky breakdown, both CG and CR fail to converge by definition, which is an uninteresting scenario. The more compelling question is to explore their convergence behavior in the absence of an unlucky breakdown throughout the iterations. Therefore, to align with our goal of investigating the convergence of CG and CR for inconsistent indefinite systems, we exclude the former case and focus on the latter, justifying the introduction of Assumption 1. Furthermore, we note that Assumption 1 is only required for indefinite settings. This is because it turns out that for positive semi-definite matrices, the only breakdown is a lucky one, which only happens at the very last iteration, and signals that the algorithm has indeed reached the grade; see Lemmas 1 and 3. In addition, in most practical applications, unlucky breakdowns are virtually never encountered, which is consistent with the observations in [42]. Therefore, we argue that Assumption 1 is reasonable in our context.
We note in passing that, in some applications, encountering a zero-curvature direction while solving 1 can be algorithmically beneficial. For instance, in Newton-type optimization methods, where $\mathbf{A}$ and $\mathbf{b}$ typically represent the Hessian and the negative gradient of a function, respectively, a zero-curvature direction can be leveraged to construct effective descent directions (see, e.g., [27, 24, 39]).
Remark 3.
One might suspect that with $\mathbf{x}_{0}$ if, at the first iteration, $\langle\mathbf{b},\mathbf{Ab}\rangle\neq 0$, subsequent iterations would also avoid an unlucky breakdown. However, as noted in [13], this is not necessarily the case. The following example,
$\mathbf{A}=\begin{bmatrix}1&0&0\\ 0&-1&0\\ 0&0&2\end{bmatrix}\quad\quad\mathbf{b}=\begin{bmatrix}1\\ -1/\sqrt{11}\\ 1,\end{bmatrix}$
shows that, even though $\left\langle\mathbf{b},\mathbf{Ab}\right\rangle\neq 0$, at the second iteration of CG, we have $\left\langle\mathbf{p}_{1},\mathbf{A}\mathbf{p}_{1}\right\rangle=0$
To end this section, we briefly outline some important, and well-known, facts about the behavior of CD methods under consistent systems. The proofs are either omitted or are deferred to Appendix B for completeness.
Fact 1 (Pseudo-inverse Solution Under Consistent Systems).
Let $\mathbf{b}\in\text{Range}(\mathbf{A})$. Under Assumption 1, Algorithms 1 and 2 converge to the pseudo-inverse solution $\mathbf{x}_{g}=\mathbf{A}^{\dagger}\mathbf{b}$ and $\mathbf{r}_{g}=\mathbf{0}$.
Fact 2.
Let $\mathbf{b}\in\text{Range}(\mathbf{A})$. Under Assumption 1, for Algorithms 1 and 2, we have $\mathbf{p}_{k}=\mathbf{0}$ if and only if $k=g$.
For consistent systems, under Assumption 1, both the residual vector (Fact 1), and the update direction (Fact 2) are zero at termination for the CD methods. Fact 1 has been shown in the works of Hayami [19, 18] for semi-definite matrices. It is a straightforward extension to the indefinite matrices with Assumption 1 in place, and the proof of Fact 1 is a direct consequence of Lemma 15, which can be found in Appendix B. Fact 2 indicates that, under consistent systems, not only does $\mathbf{r}_{g}=\mathbf{0}$ at termination, but $\mathbf{p}_{g}=\mathbf{0}$ as well. While this fact may seem trivial, it is worth noting as it sharply contrasts with the behavior of CG under inconsistent systems, where $\mathbf{Ar}$ is never zero (Theorem 1), but $\mathbf{Ap}$ is zero at termination (Lemma 1).
3 Inconsistent Systems
Under consistent systems, and in the absence of round-off errors, the orthogonality and $\mathbf{A}$-conjugacy are the distinctive features of the CD methods. These properties play a key role in the development of the characteristically clean and concise algorithmic procedure of these methods. However, as long as the algorithm has not encountered an unlucky breakdown, these properties continue to hold even in inconsistent settings. The proof of Property 1 follows the standard line of reasoning in the consistent case, and hence is deferred to Appendix A; see Lemmas 10 and 9.
Property 1.
Consider any real symmetric matrix $\mathbf{A}$, let $\{\mathbf{p}_{0},\cdots,\mathbf{p}_{k-1}\}$ and $\{\mathbf{r}_{0},\cdots,\mathbf{r}_{k-1}\}$ be the update directions and residuals generated by Algorithms 1 and 2 up to iteration $k-1$.
  1. (a)
    For both Algorithms 1 and 2, we have $\dim(\mathcal{K}_{k}(\mathbf{A},\mathbf{b}))=k$ and,
    $\text{Span}(\mathbf{r}_{0},\cdots,\mathbf{r}_{k-1})$ $=\text{Span}(\mathbf{p}_{0},\cdots,\mathbf{p}_{k-1})=\mathcal{K}_{k}(\mathbf{A},\mathbf{b})$
  2. (b)
    In CG, Algorithm 1, for all $10\leq i<j\leq k-$ 1, we have
    $\left\langle\mathbf{r}_{i},\mathbf{r}_{j}\right\rangle=0,\quad\left\langle\mathbf{p}_{i},\mathbf{Ap}_{j}\right\rangle=0\quad\text{and}\quad\left\langle\mathbf{r}_{j},\mathbf{p}_{i}\right\rangle=0;$
    while for all $10\leq j\leq i\leq k-$ 1, we have $\left\langle\mathbf{r}_{j},\mathbf{p}_{i}\right\rangle=\left\|\mathbf{r}_{i}\right\|^{2}$.
  3. (c)
    In CR, Algorithm 2, for all $10\leq i<j\leq k-$ 1, we have
    $\left\langle\mathbf{r}_{i},\mathbf{Ar}_{j}\right\rangle=0,\quad\text{and}\quad\left\langle\mathbf{Ap}_{i},\mathbf{Ap}_{j}\right\rangle=0.$
Note that Property 1 does not require Assumption 1 since these properties hold prior to a potential occurrence of an unlucky breakdown. Indeed, Assumption 1 is only required when we need to ensure that Algorithms 1 and 2 can successfully generate all iterates up to termination.
3.1 Conjugate Gradient
Under an inconsistent system, there is a non-trivial, and a prior unknown, lower bound on the norm of the residual. Indeed, $\left\|\mathbf{r}\right\|^{2}=\left\|\mathbf{b}-\mathbf{Ax}\right\|^{2}=\|\mathbf{b}^{\perp}\|^{2}+\|\tilde{\mathbf{b}}-\mathbf{Ax}\|^{2}\geq\|\mathbf{b}^{\perp}\|^{2}$. Thus, termination criterion based on $\left\|\mathbf{r}\right\|$ is no longer practical. The alternative is to observe $\left\|\mathbf{Ar}\right\|\rightarrow 0$, as $\|\mathbf{Ar}\|=0$ indicates achieving a normal solution. However, we now show that CG is fundamentally unable to recover normal solutions in such settings. While it has been empirically observed that CG exhibits numerical instability under inconsistent systems, Theorem 1 provides a theoretical analysis of CG’s failure to converge in such settings.
Theorem 1 (CG’s Inability to Recover Normal Solutions).
Let $\mathbf{b}\notin\text{Range}(\mathbf{A})$ and $\mathbf{b}\notin\text{Null}(\mathbf{A})$. Under Assumption 1, in Algorithm 1, we have $\mathbf{Ar}_{k}\neq\mathbf{0}$, for all $10\leq k\leq g-$ 1.
Proof.
We prove the claim by contradiction. Clearly, by the assumption on $\mathbf{b}$, $\mathbf{Ar}_{0}=\mathbf{Ab}\neq\mathbf{0}$. Suppose for some $11\leq k\leq g-$ 1, $\mathbf{A}\mathbf{r}_{k}=\mathbf{0}$. Using $\mathbf{p}_{k}=\mathbf{r}_{k}+\beta_{k-1}\mathbf{p}_{k-1}$, we get $\mathbf{A}\mathbf{p}_{k}=\beta_{k-1}\mathbf{A}\mathbf{p}_{k-1}$. By the assumption on $\mathbf{b}$, we have $\beta_{k-1}\neq 0$. Subsequently, from $\mathbf{r}_{k}=\mathbf{r}_{k-1}-\alpha_{k-1}\mathbf{Ap}_{k-1}$, we get $\mathbf{r}_{k}=\mathbf{r}_{k-1}-\alpha_{k-1}\mathbf{Ap}_{k}/\beta_{k-1}$, and by the inner-product with $\mathbf{p}_{k-1}$ and using Property 1-(b), it gives $\left\langle\mathbf{p}_{k-1},\mathbf{r}_{k-1}\right\rangle=0$. However, by Property 1-(b), we also have $\left\langle\mathbf{p}_{k-1},\mathbf{r}_{k-1}\right\rangle=\left\|\mathbf{r}_{k-1}\right\|^{2}\neq 0$, for $10\leq k\leq g-$ 1, which leads to a contradiction. ∎
Remark 4.
CG’s inability to recover a solution of the normal equation can be better understood through the framework from which it is derived. Specifically, CG is a particular algorithmic implementation of the Orthogonal-Residual (OR) framework [15, 4], which applies to square, but not necessarily symmetric, matrices. When restricted to PD matrices, the OR framework can be algorithmically implemented as Algorithm 1. However, in Appendix C, we demonstrate that any algorithm derived from the OR framework cannot converge to a normal solution for symmetric but inconsistent systems (see Theorem 6). Since CG is derived from OR, it inherits this “non-convergent” property.
While Theorem 1 for inconsistent systems stands in sharp contrast to Fact 1 for consistent systems, it turns out that the last iterate of CG for inconsistent settings can still be characterized, in an analogous way to Fact 2 for consistent systems.
Lemma 1.
Let $\mathbf{b}\notin\text{Range}(\mathbf{A})$. Under Assumption 1, in Algorithm 1 we have $\mathbf{Ap}_{k-1}=\mathbf{0}$ if and only if $k=g$. Furthermore, $\mathbf{p}_{g-1}\neq\mathbf{0}$.
Proof.
Let $k=g$. By Property 1-(a) and Definition 2, we note that $\mathbf{Ap}_{g-1}\in\mathcal{K}_{g}(\mathbf{A},\mathbf{b})$ and
$\mathbf{Ap}_{g-1}$ $=\sum_{j=1}^{g}c_{j-1}\mathbf{r}_{j-1}.$
In addition, from Property 1-(b), we have $\left\langle\mathbf{p}_{i-1},\mathbf{Ap}_{g-1}\right\rangle=0$, for all $11\leq i<g-$ 1. For $i=1$, $0=\left\langle\mathbf{p}_{0},\mathbf{Ap}_{g-1}\right\rangle=c_{0}\left\langle\mathbf{p}_{0},\mathbf{r}_{0}\right\rangle=c_{0}\|\mathbf{r}_{0}\|^{2}$, which since $\|\mathbf{r}_{0}\|\neq 0$ implies $c_{0}=0$. Similarly, for $12\leq i<g-$ 1,
00 0 $=\left\langle\mathbf{p}_{i-1},\mathbf{Ap}_{g-1}\right\rangle=\left\langle\mathbf{r}_{i-1}+\beta_{i-2}\mathbf{p}_{i-2},\mathbf{Ap}_{g-1}\right\rangle=\left\langle\mathbf{r}_{i-1},\mathbf{Ap}_{g-1}\right\rangle$
$=\sum_{j=1}^{g}c_{j-1}\left\langle\mathbf{r}_{i-1},\mathbf{r}_{j-1}\right\rangle=c_{i-1}\|\mathbf{r}_{i-1}\|^{2},$
where the last equality follows from Property 1-(b). Since $\|\mathbf{r}_{i-1}\|\neq 0$, it follows that $c_{i-1}=0$. As a result, we must have $\mathbf{Ap}_{g-1}=c_{g-1}\mathbf{r}_{g-1}$. Now, if $c_{g-1}=0$, then we are done. Otherwise, assuming $\mathbf{Ap}_{g-1}\not=\mathbf{0}$, since $\mathbf{r}_{g-1}\neq\mathbf{0}$, we can, in principle, proceed to the next iteration by calculating $\alpha_{g-1}=\|\mathbf{r}_{g-1}\|^{2}/\left\langle\mathbf{p}_{g-1},\mathbf{Ap}_{g-1}\right\rangle$. We first note that by Property 1-(b), we have
$c_{g-1}=\frac{\left\langle\mathbf{p}_{g-1},\mathbf{Ap}_{g-1}\right\rangle}{\left\langle\mathbf{p}_{g-1},\mathbf{r}_{g-1}\right\rangle}=\frac{\left\langle\mathbf{p}_{g-1},\mathbf{Ap}_{g-1}\right\rangle}{\left\langle\mathbf{r}_{g-1}+\beta_{g-2}\mathbf{p}_{g-2},\mathbf{r}_{g-1}\right\rangle}=\frac{\left\langle\mathbf{p}_{g-1},\mathbf{Ap}_{g-1}\right\rangle}{\|\mathbf{r}_{g-1}\|^{2}}=\frac{1}{\alpha_{g-1}}.$
It follows that $\mathbf{r}_{g}=\mathbf{r}_{g-1}-\alpha_{g-1}\mathbf{Ap}_{g-1}=\mathbf{r}_{g-1}-\alpha_{g-1}c_{g-1}\mathbf{r}_{g-1}=\mathbf{0}$, which is impossible since $\mathbf{b}\notin\text{Range}(\mathbf{A})$. By reductio ad absurdum, this implies $\mathbf{Ap}_{g-1}=\mathbf{0}$.
Conversely, let $\mathbf{Ap}_{k-1}=\mathbf{0}$ for some $1\leq k\leq g$. Firstly, we show that $\mathbf{p}_{k-1}\not=\mathbf{0}$. Clearly, if $k=1$, we have $\mathbf{p}_{0}=\mathbf{b}\neq\mathbf{0}$. If $k\geq 2$, from $\mathbf{p}_{k-1}=\mathbf{r}_{k-1}+\beta_{k-2}\mathbf{p}_{k-2}$ and Property 1-(b), it follows that
$\left\|\mathbf{p}_{k-1}\right\|^{2}$ $=\left\|\mathbf{r}_{k-1}\right\|^{2}+2\beta_{k-2}\left\langle\mathbf{r}_{k-1},\mathbf{p}_{k-2}\right\rangle+\beta_{k-2}^{2}\left\|\mathbf{p}_{k-2}\right\|^{2}$
$=\left\|\mathbf{r}_{k-1}\right\|^{2}+\beta_{k-2}^{2}\left\|\mathbf{p}_{k-2}\right\|^{2}\geq\left\|\mathbf{r}_{k-1}\right\|^{2}\neq 0.$
As a result, since $\mathbf{p}_{k-1}\in\mathcal{K}_{k}(\mathbf{A},\mathbf{b})$ and $\mathbf{p}_{k-1}\neq\mathbf{0}$, we can write $\mathbf{p}_{k-1}=\sum_{i=0}^{k-1}c_{i}\mathbf{A}^{i}\mathbf{b}$ and at least one $c_{j}$ must be non-zero. Hence, $\mathbf{0}=\mathbf{Ap}_{k-1}=\sum_{i=0}^{k-1}c_{i}\mathbf{A}^{i+1}\mathbf{b}$, which implies $\dim(\mathcal{K}_{k+1}(\mathbf{A},\mathbf{b}))=\dim(\mathcal{K}_{k}(\mathbf{A},\mathbf{b}))$, and so we must have $k=g$. ∎
Remark 5.
There is a subtlety in how the grade of $\mathbf{b}$ with respect to $\mathbf{A}$ is reflected in the total number of iterations before termination. If $\mathbf{b}\in\text{Range}(\mathbf{A})$, CG terminates when $k=g$ (Facts 1 and 2), while for $\mathbf{b}\notin\text{Range}(\mathbf{A})$, the termination occurs at $k=g-1$ (Lemma 1). However, in both cases, the total number of iterations is determined by the number of non-zero $\mathbf{b}$-relevant eigenvalues [^7] These are distinct eigenvalues whose corresponding eigenspace is not entirely orthogonal to $\mathbf{b}$; see the work [27] for more details of $\mathbf{A}$. Indeed, suppose $\mathbf{b}\notin\text{Range}(\mathbf{A})$ and consider its decomposition as in 2. Let $g$ and $\tilde{g}$ denote the grade of $\mathbf{b}$ and $\tilde{\mathbf{b}}$, respectively, with respect to $\mathbf{A}$. CG$(\mathbf{A},\tilde{\mathbf{b}})$ and CG$(\mathbf{A},\mathbf{b})$ terminate after $\tilde{g}$ and $g-1$ iterations, respectively. On the other hand, noting that $\tilde{\mathbf{b}}$ is orthogonal to the eigenspace corresponding to the zero eigenvalues of $\mathbf{A}$, for any eigenvector $\mathbf{v}$ of $\mathbf{A}$, we have
$\left\langle\mathbf{b},\mathbf{v}\right\rangle=\langle\tilde{\mathbf{b}},\mathbf{v}\rangle+\langle\mathbf{b}^{\perp},\mathbf{v}\rangle=\left\{\begin{array}[]{ll}\langle\tilde{\mathbf{b}},\mathbf{v}\rangle,&\text{if }\mathbf{v}\in\text{Range}(\mathbf{A}),\\ \\ \langle\mathbf{b}^{\perp},\mathbf{v}\rangle,&\text{if }\mathbf{v}\in\text{Null}(\mathbf{A}).\end{array}\right.$
That is, the number of $\mathbf{b}$-relevant eigenvalues is exactly one more than the number of $\tilde{\mathbf{b}}$-relevant ones. As a result, from Lemma 1 in the work of Liu and Roosta[27], it follows that $\tilde{g}=g-1$, which implies that both situations result in precisely the same number of iterations.
Theorem 1 implies that, for inconsistent systems, CG’s output might not necessarily correspond to any solutions. Once $k=g-1$, by Lemma 1, $\alpha_{g-1}$ is no longer well-defined and the update direction $\mathbf{p}_{g-1}$ is not used. One may naïvely attempt to construct $\mathbf{x}_{g}=\mathbf{x}_{g-1}+\alpha\mathbf{p}_{g-1}$ for some suitable $\alpha$ in an effort to recover a normal solution. However, for any such update $\mathbf{x}_{g}=\mathbf{x}_{g-1}+\alpha\mathbf{p}_{g-1}$, we have $\mathbf{r}_{g}=\mathbf{r}_{g-1}$, and hence by Theorem 1, $\mathbf{x}_{g}$ cannot be a normal solution. Nonetheless, we now show that a modification to CG theoretically allows for the recovery of the pseudo-inverse solution in inconsistent settings.
To do this, we make two key observations. First, by Lemma 15 in Appendix B, the Krylov subspace $\mathcal{K}_{g}(\mathbf{A},\mathbf{b})$ contains a normal solution. Denoting this normal solution as $\mathbf{x}^{\star}$, under Assumption 1 and using Property 1-(a), we can write
$\mathbf{x}^{\star}=\sum_{k=1}^{g}\eta_{k-1}\mathbf{p}_{k-1},$ (6)
and the aim is thus to find the suitable scalars $\eta_{i}$’s. Next, if the vector $\mathbf{b}^{\perp}$ can be somehow calculated as part of CG iterations, then it might be possible to eliminate its contribution in a suitable way to recover the pseudo-inverse solution. Fortunately, it turns out that $\mathbf{b}^{\perp}$ is readily available at the last iteration of CG.
Lemma 2.
Let $\mathbf{b}\notin\text{Range}(\mathbf{A})$. Under Assumption 1, in Algorithm 1, we have
$\mathbf{p}_{g-1}=\frac{\|\mathbf{p}_{g-1}\|^{2}}{\|\mathbf{r}_{g-1}\|^{2}}\mathbf{b}^{\perp}.$
Proof.
By Lemma 12 in Appendix B, we know that $(\mathbf{I}-\mathbf{A}^{\dagger}\mathbf{A})\mathbf{p}_{g-1}=a_{g-1}\mathbf{b}^{\perp}$, which by Lemma 1 gives $\mathbf{p}_{g-1}=a_{g-1}\mathbf{b}^{\perp}$. Furthermore, Lemma 1 gives $\left\langle\mathbf{p}_{g-1},\mathbf{b}\right\rangle=\left\langle\mathbf{p}_{g-1},\tilde{\mathbf{b}}\right\rangle+\left\langle\mathbf{p}_{g-1},\mathbf{b}^{\perp}\right\rangle=\left\langle\mathbf{p}_{g-1},\mathbf{b}^{\perp}\right\rangle$. Hence,
$a_{g-1}=\frac{\left\|\mathbf{p}_{g-1}\right\|^{2}}{\left\langle\mathbf{p}_{g-1},\mathbf{b}^{\perp}\right\rangle}=\frac{\left\|\mathbf{p}_{g-1}\right\|^{2}}{\left\langle\mathbf{p}_{g-1},\mathbf{b}\right\rangle}.$
Finally, using Property 1-(b) and the definition of $\beta_{k-1}$, we get $\left\langle\mathbf{p}_{g-1},\mathbf{b}\right\rangle=\|\mathbf{r}_{g-1}\|^{2}$. ∎
We now express 6 as
$\mathbf{x}^{\star}=\sum_{k=1}^{g}\eta_{k-1}\mathbf{p}_{k-1}=\sum_{k=1}^{g}(\alpha_{k-1}+\gamma_{k-1})\mathbf{p}_{k-1},$ (7)
where $\alpha_{k}$’s are from Algorithm 1 and $\gamma_{k}$’s are unknown constant that we aim to find. It follows that
$\frac{\|\mathbf{r}_{g-1}\|^{2}}{\|\mathbf{p}_{g-1}\|^{2}}\mathbf{p}_{g-1}$ $=\mathbf{b}^{\perp}=\mathbf{b}-\mathbf{A}\mathbf{x}^{\star}=\mathbf{b}-\sum_{k=1}^{g}\alpha_{k-1}\mathbf{Ap}_{k-1}-\sum_{k=1}^{g}\gamma_{k-1}\mathbf{Ap}_{k-1}=\mathbf{r}_{g-1}-\sum_{k=1}^{g-1}\gamma_{k-1}\mathbf{Ap}_{k-1},$
where the first equality uses Lemma 2, and the fourth equality employs Lemma 1 and the fact that in Algorithm 1, $\mathbf{r}_{k}=\mathbf{r}_{k-1}-\alpha_{k-1}\mathbf{Ap}_{k-1}$. Now, for each $11\leq k\leq g-$ 1, we have
$\frac{\|\mathbf{r}_{g-1}\|^{2}}{\|\mathbf{p}_{g-1}\|^{2}}\left\langle\mathbf{p}_{k-1},\mathbf{p}_{g-1}\right\rangle$ $=\left\langle\mathbf{p}_{k-1},\mathbf{r}_{g-1}\right\rangle-\gamma_{k-1}\left\langle\mathbf{p}_{k-1},\mathbf{Ap}_{k-1}\right\rangle,$
which using Property 1-(b) gives
$\gamma_{k-1}$ $=-\frac{\|\mathbf{r}_{g-1}\|^{2}\left\langle\mathbf{p}_{k-1},\mathbf{p}_{g-1}\right\rangle}{\|\mathbf{p}_{g-1}\|^{2}\left\langle\mathbf{p}_{k-1},\mathbf{Ap}_{k-1}\right\rangle}.$
Furthermore,
$\left\langle\mathbf{p}_{k-1},\mathbf{p}_{g-1}\right\rangle$ $=\left\langle\mathbf{p}_{k-1},\mathbf{r}_{g-1}+\beta_{g-2}\mathbf{p}_{g-2}\right\rangle=\beta_{g-2}\left\langle\mathbf{p}_{k-1},\mathbf{p}_{g-2}\right\rangle=\beta_{g-2}\beta_{g-3}\left\langle\mathbf{p}_{k-1},\mathbf{p}_{g-3}\right\rangle=\ldots$
$=\prod_{j=2}^{g-(k-1)}\beta_{g-j}\|\mathbf{p}_{k-1}\|^{2}=\frac{\|\mathbf{r}_{g-1}\|^{2}}{\|\mathbf{r}_{k-1}\|^{2}}\|\mathbf{p}_{k-1}\|^{2},$
which gives,
$\gamma_{k-1}=-\frac{\|\mathbf{r}_{g-1}\|^{4}\|\mathbf{p}_{k-1}\|^{2}}{\|\mathbf{p}_{g-1}\|^{2}\|\mathbf{r}_{k-1}\|^{2}\left\langle\mathbf{p}_{k-1},\mathbf{Ap}_{k-1}\right\rangle},$
for $11\leq k\leq g-$ 1. Using these scalars, coupled with the particular choice $\gamma_{g-1}=0$, $\mathbf{x}^{\star}$ in 7 now amounts to a normal solution. To find the pseudo-inverse solution, we note that, Lemma 12 implies
$(\mathbf{I}-\mathbf{A}^{\dagger}\mathbf{A})\mathbf{x}^{\star}=(\mathbf{I}-\mathbf{A}^{\dagger}\mathbf{A})\sum_{k=1}^{g}\eta_{k-1}\mathbf{p}_{k-1}=\left(\sum_{k=1}^{g}\eta_{k-1}a_{k-1}\right)\mathbf{b}^{\perp}=c\mathbf{p}_{g-1},$
where $c=\sum_{k=1}^{g}\eta_{k-1}a_{k-1}$. However, using Lemma 1,
$c=\frac{\left\langle\mathbf{p}_{g-1},\mathbf{x}^{\star}\right\rangle}{\|\mathbf{p}_{g-1}\|^{2}}.$
Hence, $(\mathbf{I}-\mathbf{A}^{\dagger}\mathbf{A})\mathbf{x}^{\star}=\left\langle\mathbf{p}_{g-1},\mathbf{x}^{\star}\right\rangle\mathbf{p}_{g-1}/\|\mathbf{p}_{g-1}\|^{2}$, which finally gives
$\mathbf{x}^{\star}-\frac{\left\langle\mathbf{p}_{g-1},\mathbf{x}^{\star}\right\rangle}{\|\mathbf{p}_{g-1}\|^{2}}\mathbf{p}_{g-1}$ $=\mathbf{A}^{\dagger}\mathbf{A}\mathbf{x}^{\star}=\mathbf{A}^{\dagger}\mathbf{b}.$
Note that the left-hand side can also be regarded as the orthogonal projection of $\mathbf{x}^{\star}$ onto the orthogonal complement of $\text{Range}(\mathbf{p}_{g-1})$. This leads us to Algorithms 3 and 2.
Algorithm 3 Conjugate Gradient with Pseudo-inverse Solution
1:$\mathbf{A}$ and $\mathbf{b}$
2:Let $k=1$, $\mathbf{r}_{0}=\mathbf{p}_{0}=\mathbf{b}$
3:while $\|\mathbf{Ap}_{k-1}\|\neq 0$ do $\triangleright$ Termination Condition
4:     $\alpha_{k-1}=\langle\mathbf{r}_{k-1},\mathbf{r}_{k-1}\rangle/\left\langle\mathbf{p}_{k-1},\mathbf{Ap}_{k-1}\right\rangle$triangleright Classical CG: Steps 3 - 7
5:     $\mathbf{x}_{k}=\mathbf{x}_{k-1}+\alpha_{k-1}\mathbf{p}_{k-1}$
6:     $\mathbf{r}_{k}=\mathbf{r}_{k-1}-\alpha_{k-1}\mathbf{Ap}_{k-1}$
7:     $\beta_{k-1}=\left\langle\mathbf{r}_{k},\mathbf{r}_{k}\right\rangle/\left\langle\mathbf{r}_{k-1},\mathbf{r}_{k-1}\right\rangle$
8:     $\mathbf{p}_{k}=\mathbf{r}_{k}+\beta_{k-1}\mathbf{p}_{k-1}$
9:     $\gamma_{k-1}=\left\langle\mathbf{p}_{k-1},\mathbf{p}_{k-1}\right\rangle/(\left\langle\mathbf{r}_{k-1},\mathbf{r}_{k-1}\right\rangle\left\langle\mathbf{p}_{k-1},\mathbf{Ap}_{k-1}\right\rangle)$triangleright Corrective Procedure: Steps 8 - 9
10:     $\mathbf{x}^{\prime}_{k}=\mathbf{x}_{k-1}^{\prime}+\gamma_{k-1}\mathbf{p}_{k-1}$
11:     $k=k+1$
12:end while
13:if $\|\mathbf{p}_{k-1}\|\neq\mathbf{0}$ then $\triangleright$ i.e., Inconsistent System
14:     $\mathbf{x}^{\star}=\mathbf{x}_{k-1}-(\|\mathbf{r}_{k-1}\|^{4}/\|\mathbf{p}_{k-1}\|^{2})\mathbf{x}_{k-1}^{\prime}$triangleright Normal Solution
15:     $\mathbf{x}^{+}=\mathbf{x}^{\star}-(\left\langle\mathbf{p}_{k-1},\mathbf{x}^{\star}\right\rangle/\|\mathbf{p}_{k-1}\|^{2})\mathbf{p}_{k-1}$triangleright Pseudo-inverse Solution
16$\triangleright$ i.e., Consistent System
17:     $\mathbf{x}^{+}=\mathbf{x}_{k-1}$
18:end if
19:return $\mathbf{x}^{+}$
Theorem 2 (Pseudo-inverse Solution with Algorithm 3).
Under Assumption 1, in Algorithm 3 we have $\mathbf{x}^{+}=\mathbf{A}^{\dagger}\mathbf{b}$.
Remark 6.
We highlight that while, theoretically, Algorithm 3 provides a means to extract the pseudo-inverse solution at termination, the extreme numerical instability of CG for inconsistent systems causes the theory to fall short of practical observations. The instability of CG under inconsistent systems is widely observed in many applications and also documented in the works [23, 35]; see also the experiments in Section 4.
3.2 Conjugate Residual
In this section, we show that under Assumption 1, CR is able to converge to a normal solution for inconsistent systems (Theorem 3). In doing so, we will also demonstrate the similarities between CR and CG, while highlighting additional advantages of CR. Similar to CG, we will then introduce a few additional steps within CR to recover the pseudo-inverse solution (Theorem 4). Finally, we will show that, as long as an unlucky breakdown has not occurred, CR generates iterates that are identical to those of MINRES for any symmetric matrix $\mathbf{A}$ and any vector $\mathbf{b}$ (Theorem 5).
3.2.1 Convergence Behavior
The convergent behavior of CR under various conditions have been extensively studied (see Table 2). Originally developed by Hestenes and Stiefel [22], CR was designed for positive definite matrices and can be straightforwardly applied to negative definite matrices. Luenberger [30] later extended CR to indefinite, albeit non-singular, matrices. For inconsistent systems, Hayami and Sugihara [18, 21, 20] established that for inconsistent systems, the necessary and sufficient condition for CR to avoid unlucky breakdown is that the matrix $\mathbf{A}$ is positive semi-definite. In particular, they demonstrated that for any indefinite $\mathbf{A}$, there exists a right-hand side vector $\mathbf{b}$ that will trigger an unlucky breakdown. While this is a crucial insight, it remains of interest to investigate the behavior of CR for the many other vectors $\mathbf{b}$ that do not cause this breakdown. We do that here. Specifically, we show that for any symmetric matrix $\mathbf{A}$ and any vector $\mathbf{b}$, if no unlucky breakdown occurs, CR will indeed converge to a normal solution. Furthermore, we devise a few extra steps for CR to obtain the pseudo-inverse solution.
Consistent: $\mathbf{b}\in\text{Range}(\mathbf{A})$ Inconsistent: $\mathbf{b}\notin\text{Range}(\mathbf{A})$
(Semi-)Definite Hestenes & Stiefel [22] Hayami & Sugihara [18, 21, 20]
Indefinite Luenberger [30] & Greenbaum [15] Theorems 3 and 4
Table 2: Relevant studies on the convergent behavior of CR under various conditions.
Remark 7.
There are many algorithms considering the alternative Krylov subspaces, like $\mathcal{K}(\mathbf{A},\mathbf{Ab})$ or $\mathcal{K}(\mathbf{A^{\top}A},\mathbf{Ab})$, instead of $\mathcal{K}(\mathbf{A},\mathbf{b})$, for example MR-II / RRMR [16] and variants of CG and CR [2, 11]. The advantages of these alternative Krylov subspaces are that the vector $\mathbf{Ab}$ is in the range of $\mathbf{A}$ and, hence, avoids unnecessary complications arising from the inconsistent systems. However, it is also empirically shown that these alternative subspaces may not produce better results than $\mathcal{K}(\mathbf{A},\mathbf{b})$ and has more computational costs, see the work [3] discussing MINRES and RRMR. It is also shown that the iterates of Krylov subspace methods defined in $\mathcal{K}(\mathbf{A},\mathbf{Ab})$ tend to have a slower convergence than $\mathcal{K}(\mathbf{A},\mathbf{b})$, see the work[7] comparing SYMMLQ and CG.
One of the similarities between CR and CG is that, as in Lemma 1 for CG, at termination for CR we also have $\mathbf{p}_{g-1}\in\text{Null}(\mathbf{A})$.
Lemma 3.
Let $\mathbf{b}\notin\text{Range}(\mathbf{A})$. Under Assumption 1, in Algorithm 2, we have $\mathbf{Ap}_{k-1}=\mathbf{0}$ if and only if $k=g$. Furthermore, $\mathbf{p}_{g-1}\neq\mathbf{0}$.
Proof.
The proof follows a similar line of reasoning as that of Lemma 1. Let $k=g$. If $g=1$, since $\mathbf{b}\notin\text{Range}(\mathbf{A})$, we must have $\mathbf{p}_{0}=\mathbf{b}\in\text{Null}(\mathbf{A})$ and the claim holds. So consider $g>1$. By Property 1-(a) and Definition 2, we note that $\mathbf{Ap}_{g-1}\in\mathcal{K}_{g}(\mathbf{A},\mathbf{b})$ and
$\mathbf{Ap}_{g-1}$ $=\sum_{j=1}^{g}c_{j-1}\mathbf{r}_{j-1}.$
In addition, from Property 1-(c), we have $\left\langle\mathbf{A}\mathbf{p}_{i-1},\mathbf{Ap}_{g-1}\right\rangle=0$, for all $11\leq i<g-$ 1. For $i=1$, $0=\left\langle\mathbf{A}\mathbf{p}_{0},\mathbf{Ap}_{g-1}\right\rangle=c_{0}\left\langle\mathbf{A}\mathbf{p}_{0},\mathbf{r}_{0}\right\rangle=c_{0}\left\langle\mathbf{A}\mathbf{r}_{0},\mathbf{r}_{0}\right\rangle$. We note that $\left\langle\mathbf{A}\mathbf{r}_{0},\mathbf{r}_{0}\right\rangle\neq 0$ by Assumption 1. Hence, we must have $c_{0}=0$. Similarly, for $12\leq i<g-$ 1,
00 0 $=\left\langle\mathbf{A}\mathbf{p}_{i-1},\mathbf{Ap}_{g-1}\right\rangle=\left\langle\mathbf{A}\mathbf{r}_{i-1}+\beta_{i-2}\mathbf{A}\mathbf{p}_{i-2},\mathbf{Ap}_{g-1}\right\rangle=\left\langle\mathbf{A}\mathbf{r}_{i-1},\mathbf{Ap}_{g-1}\right\rangle$
$=\sum_{j=1}^{g}c_{j-1}\left\langle\mathbf{A}\mathbf{r}_{i-1},\mathbf{r}_{j-1}\right\rangle=c_{i-1}\left\langle\mathbf{A}\mathbf{r}_{i-1},\mathbf{r}_{i-1}\right\rangle,$
where the last equality follows from Property 1-(c). Again, since $\left\langle\mathbf{A}\mathbf{r}_{i-1},\mathbf{r}_{i-1}\right\rangle\neq 0$, by Assumption 1, it follows that $c_{i-1}=0$. As a result, we must have got $\mathbf{Ap}_{g-1}=c_{g-1}\mathbf{r}_{g-1}$. Now, if $c_{g-1}=0$, then we are done. Otherwise, assuming $\mathbf{Ap}_{g-1}\not=\mathbf{0}$, since $\mathbf{r}_{g-1}\neq\mathbf{0}$, the constant $\alpha_{g-1}=\left\langle\mathbf{r}_{g-1},\mathbf{A}\mathbf{r}_{g-1}\right\rangle/\left\langle\mathbf{A}\mathbf{p}_{g-1},\mathbf{Ap}_{g-1}\right\rangle$ can be, in principle, calculated in the next iterate. We first note that by Property 1-(c), we have
$c_{g-1}$ $=\frac{\left\langle\mathbf{A}\mathbf{p}_{g-1},\mathbf{Ap}_{g-1}\right\rangle}{\left\langle\mathbf{A}\mathbf{p}_{g-1},\mathbf{r}_{g-1}\right\rangle}=\frac{\left\langle\mathbf{A}\mathbf{p}_{g-1},\mathbf{Ap}_{g-1}\right\rangle}{\left\langle\mathbf{A}\mathbf{r}_{g-1}+\beta_{g-2}\mathbf{A}\mathbf{p}_{g-2},\mathbf{r}_{g-1}\right\rangle}$
$=\frac{\left\langle\mathbf{A}\mathbf{p}_{g-1},\mathbf{Ap}_{g-1}\right\rangle}{\left\langle\mathbf{r}_{g-1},\mathbf{Ar}_{g-1}\right\rangle}=\frac{1}{\alpha_{g-1}}.$
It follows that $\mathbf{r}_{g}=\mathbf{r}_{g-1}-\alpha_{g-1}\mathbf{Ap}_{g-1}=\mathbf{r}_{g-1}-\alpha_{g-1}c_{g-1}\mathbf{r}_{g-1}=\mathbf{0}$, which is impossible since $\mathbf{b}\notin\text{Range}(\mathbf{A})$. By reductio ad absurdum, this implies $\mathbf{Ap}_{g-1}=\mathbf{0}$.
Conversely, let $1\leq k\leq g$ be the first iteration where $\mathbf{Ap}_{k-1}=\mathbf{0}$. We first show that $\mathbf{p}_{k-1}\neq\mathbf{0}$. If $k=1$, we have $\mathbf{p}_{0}=\mathbf{b}\neq\mathbf{0}$. Suppose $k\geq 2$ and $\mathbf{p}_{k-1}=\mathbf{0}$. Clearly, $\beta_{k-2}\neq 0$ as otherwise $\mathbf{0}=\mathbf{p}_{k-1}=\mathbf{r}_{k-1}+\beta_{k-2}\mathbf{p}_{k-2}$ also implies $\mathbf{r}_{k-1}=\mathbf{0}$ which by assumption on $\mathbf{b}$ is impossible. However, by Property 1-(a), $\mathbf{r}_{k-1}$ and $\mathbf{p}_{k-2}$ must be linearly independent and we arrive at a contradiction. As a result, since $\mathbf{p}_{k-1}\in\mathcal{K}_{k}(\mathbf{A},\mathbf{b})$ and $\mathbf{p}_{k-1}\neq\mathbf{0}$, we can write $\mathbf{p}_{k-1}=\sum_{i=0}^{k-1}c_{i}\mathbf{A}^{i}\mathbf{b}$ and at least one $c_{j}$ must be non-zero. Hence, $\mathbf{0}=\mathbf{Ap}_{k-1}=\sum_{i=0}^{k-1}c_{i}\mathbf{A}^{i+1}\mathbf{b}$, which implies $\dim(\mathcal{K}_{k+1}(\mathbf{A},\mathbf{b}))=\dim(\mathcal{K}_{k}(\mathbf{A},\mathbf{b}))$, and so we must have $k=g$. ∎
While Lemmas 1 and 3 highlight a similarity between CG and CR, Theorems 1 and 3 showcase a drastic difference. Specifically, we now show that, in sharp contrast to CG, CR always recovers a normal solution.
Theorem 3 (CR’s Ability to Recover Normal Solution).
Let $\mathbf{b}\notin\text{Range}(\mathbf{A})$. Under Assumption 1, in Algorithm 2 we have $\mathbf{Ar}_{k-1}=\mathbf{0}$ if and only if $k=g$.
Proof.
Suppose $k=g$. From $\mathbf{p}_{g-1}=\mathbf{r}_{g-1}+\beta_{g-2}\mathbf{p}_{g-2}$ and Lemma 3, we have $\left\langle\mathbf{r}_{g-1},\mathbf{Ar}_{g-1}\right\rangle+\beta_{g-2}\left\langle\mathbf{r}_{g-1},\mathbf{Ap}_{g-2}\right\rangle=0$, which using Property 1-(c) gives $\left\langle\mathbf{r}_{g-1},\mathbf{Ar}_{g-1}\right\rangle=0$, implying $\beta_{g-2}=0$ by its definition in Algorithm 2. Hence, $\mathbf{p}_{g-1}=\mathbf{r}_{g-1}$ and the result follows from Lemma 3.
Conversely, suppose $\mathbf{Ar}_{k-1}=\mathbf{0}$ for some iteration $1\leq k\leq g$. From $\mathbf{p}_{k-1}=\mathbf{r}_{k-1}+\beta_{k-2}\mathbf{p}_{k-2}$, we get $\mathbf{Ap}_{k-1}=\beta_{k-2}\mathbf{Ap}_{k-2}$, which using Property 1-(c) gives $\|\mathbf{Ap}_{k-1}\|^{2}=\beta_{k-2}\left\langle\mathbf{Ap}_{k-1},\mathbf{Ap}_{k-2}\right\rangle=0$. Hence, $\mathbf{Ap}_{k-1}=\mathbf{0}$, which by Lemma 3 can only happen when $k=g$. ∎
Remark 8.
For both Algorithms 1 and 2, we always have $\mathbf{p}_{0}=\mathbf{r}_{0}$. However, from the proof of Theorem 3, it can be seen that under the assumptions of Theorem 3, in Algorithm 2, we also have $\mathbf{p}_{g-1}=\mathbf{r}_{g-1}$. It is important to note that this property is specific to CR, and CG does not exhibit the same characteristics.
While, by Theorem 3, CR is guaranteed to converge to a normal solution, we next show that we can in fact obtain the pseudo-inverse solution from the last iterate of CR through a straightforward procedure. This stands in sharp contrast to CG, where a similar procedure requires non-trivial modifications to the original algorithm. To achieve this, we first establish several technical lemmas.
Lemma 4.
Consider the iterates generated by the procedures CR$(\mathbf{A},\tilde{\mathbf{b}})$ and CR$(\mathbf{A},\mathbf{b})$, and let where $g$ be the grade of $\mathbf{b}$ with respect to $\mathbf{A}$. For $1\leq k\leq g$, we have $\tilde{\alpha}_{k-1}=\alpha_{k-1}$, $\tilde{\beta}_{k-1}=\beta_{k-1}$, and $\tilde{\mathbf{x}}_{k-1}=\mathbf{A}^{\dagger}\mathbf{Ax}_{k-1}$.
Proof.
We first note that, for any $\mathbf{x}$, we have $(\mathbf{I}-\mathbf{A}^{\dagger}\mathbf{A})\mathbf{r}=\mathbf{b}^{\perp}$. To show the claim, we also need to show that $\tilde{\mathbf{p}}_{k-1}=\mathbf{A}^{\dagger}\mathbf{A}\mathbf{p}_{k-1}$, and $\tilde{\mathbf{r}}_{k-1}=\mathbf{A}^{\dagger}\mathbf{A}\mathbf{r}_{k-1}$. The proof is done by induction. For $k=1$, we have
$\mathbf{A}^{\dagger}\mathbf{A}\mathbf{r}_{0}=\mathbf{A}^{\dagger}\mathbf{A}(\tilde{\mathbf{b}}+\mathbf{b}^{\perp})=\tilde{\mathbf{b}}=\tilde{\mathbf{r}}_{0},$
$\mathbf{A}^{\dagger}\mathbf{A}\mathbf{p}_{0}=\mathbf{A}^{\dagger}\mathbf{A}(\tilde{\mathbf{b}}+\mathbf{b}^{\perp})=\tilde{\mathbf{b}}=\tilde{\mathbf{p}}_{0},$
and
$\alpha_{0}=\frac{\left\langle\mathbf{r}_{0},\mathbf{A}\mathbf{r}_{0}\right\rangle}{\|\mathbf{A}\mathbf{p}_{0}\|^{2}}=\frac{\left\langle(\tilde{\mathbf{b}}+\mathbf{b}^{\perp}),\mathbf{A}(\tilde{\mathbf{b}}+\mathbf{b}^{\perp})\right\rangle}{\|\mathbf{A}(\tilde{\mathbf{b}}+\mathbf{b}^{\perp})\|^{2}}=\frac{\left\langle\tilde{\mathbf{b}},\mathbf{A}\tilde{\mathbf{b}}\right\rangle}{\|\mathbf{A}\tilde{\mathbf{b}}\|^{2}}=\tilde{\alpha}_{0}.$
Furthermore, $\mathbf{A}^{\dagger}\mathbf{A}\mathbf{r}_{1}=\mathbf{A}^{\dagger}\mathbf{A}(\mathbf{r}_{0}-\alpha_{0}\mathbf{A}\mathbf{p}_{0})=\tilde{\mathbf{r}}_{0}-\tilde{\alpha}_{0}\mathbf{A}\tilde{\mathbf{p}}_{0}=\tilde{\mathbf{r}}_{1}$, where we have used the fact $\mathbf{A}^{\dagger}\mathbf{A}=\mathbf{AA}^{\dagger}$. So,
$\beta_{0}=\frac{\left\langle\mathbf{r}_{1},\mathbf{A}\mathbf{r}_{1}\right\rangle}{\left\langle\mathbf{r}_{0},\mathbf{A}\mathbf{r}_{0}\right\rangle}=\frac{\left\langle(\tilde{\mathbf{r}}_{1}+\mathbf{b}^{\perp}),\mathbf{A}(\tilde{\mathbf{r}}_{1}+\mathbf{b}^{\perp})\right\rangle}{\left\langle(\tilde{\mathbf{b}}+\mathbf{b}^{\perp}),\mathbf{A}(\tilde{\mathbf{b}}+\mathbf{b}^{\perp})\right\rangle}=\frac{\left\langle\tilde{\mathbf{r}}_{1}\mathbf{A}\tilde{\mathbf{r}}_{1}\right\rangle}{\left\langle\tilde{\mathbf{b}},\mathbf{A}\tilde{\mathbf{b}}\right\rangle}=\tilde{\beta}_{0}.$
Next, assume the claim holds for all $k-1$. It follows that
$\mathbf{A}^{\dagger}\mathbf{A}\mathbf{r}_{k}$ $=\mathbf{A}^{\dagger}\mathbf{A}(\mathbf{r}_{k-1}-\alpha_{k-1}\mathbf{A}\mathbf{p}_{k-1})=\tilde{\mathbf{r}}_{k-1}-\tilde{\alpha}_{k-1}\mathbf{A}\tilde{\mathbf{p}}_{k-1}=\tilde{\mathbf{r}}_{k},$
$\mathbf{A}^{\dagger}\mathbf{A}\mathbf{p}_{k}$ $=\mathbf{A}^{\dagger}\mathbf{A}(\mathbf{r}_{k}+\beta_{k-1}\mathbf{p}_{k-1})=\tilde{\mathbf{r}}_{k}+\tilde{\beta}_{k-1}\tilde{\mathbf{p}}_{k-1}=\tilde{\mathbf{p}}_{k},$
$\alpha_{k}$ $=\frac{\left\langle\mathbf{r}_{k},\mathbf{A}\mathbf{r}_{k}\right\rangle}{\|\mathbf{A}\mathbf{p}_{k}\|^{2}}=\frac{\left\langle\tilde{\mathbf{r}}_{k},\mathbf{A}\tilde{\mathbf{r}}_{k}\right\rangle}{\|\mathbf{A}\tilde{\mathbf{p}}_{k}\|^{2}}=\tilde{\alpha}_{k}.$
Also, since $\mathbf{A}^{\dagger}\mathbf{A}\mathbf{r}_{k+1}=\mathbf{A}^{\dagger}\mathbf{A}(\mathbf{r}_{k}-\alpha_{k}\mathbf{A}\mathbf{p}_{k})=\tilde{\mathbf{r}}_{k}-\tilde{\alpha}_{k}\mathbf{A}\tilde{\mathbf{p}}_{k}=\tilde{\mathbf{r}}_{k+1}$, we obtain
$\beta_{k}$ $=\frac{\left\langle\mathbf{r}_{k+1},\mathbf{A}\mathbf{r}_{k+1}\right\rangle}{\left\langle\mathbf{r}_{k},\mathbf{A}\mathbf{r}_{k}\right\rangle}=\frac{\left\langle\tilde{\mathbf{r}}_{k+1},\mathbf{A}\tilde{\mathbf{r}}_{k+1}\right\rangle}{\left\langle\tilde{\mathbf{r}}_{k},\mathbf{A}\tilde{\mathbf{r}}_{k}\right\rangle}=\tilde{\beta}_{k}.$
Finally, for any $1\leq k\leq g$, we have
$\mathbf{x}_{k}=\sum_{i=1}^{k}\alpha_{i-1}\mathbf{p}_{i-1},$
which implies
$\mathbf{A}^{\dagger}\mathbf{A}\mathbf{x}_{k}=\mathbf{A}^{\dagger}\mathbf{A}\sum_{i=1}^{k}\alpha_{i-1}\mathbf{p}_{i-1}=\sum_{i=1}^{k}\tilde{\alpha}_{i-1}\tilde{\mathbf{p}}_{i-1}=\tilde{\mathbf{x}}_{k}.$
We note that, as alluded to in Remark 5, both procedures CR$(\mathbf{A},\tilde{\mathbf{b}})$ and CR$(\mathbf{A},\mathbf{b})$ undergo precisely the same number of iterations. Hence, the assertion of Lemma 4 holds true for all iterations in both procedures.
Since $\tilde{\mathbf{b}}\in\text{Range}(\mathbf{A})$, by Fact 1, for the procedure CR($\mathbf{A},\tilde{\mathbf{b}}$), we have $\mathbf{A}^{\dagger}\mathbf{A}\mathbf{x}_{g-1}=\tilde{\mathbf{x}}_{g-1}=\tilde{\mathbf{x}}_{\tilde{g}}=\mathbf{A}^{\dagger}\tilde{\mathbf{b}}=\mathbf{A}^{\dagger}\mathbf{b}=\mathbf{x}^{+}$. Now, using a similar line of reasoning as that leading up to Theorem 2 for CG, by utilizing Lemma 12, for some constant $c$, we have
$(\mathbf{I}-\mathbf{A}^{\dagger}\mathbf{A})\mathbf{x}_{g-1}$ $=(\mathbf{I}-\mathbf{A}^{\dagger}\mathbf{A})\sum_{k=1}^{g}\alpha_{k-1}\mathbf{p}_{k-1}=c\mathbf{p}_{g-1},$
which using Lemma 3 gives $c=\left\langle\mathbf{p}_{g-1},\mathbf{x}_{g-1}\right\rangle/\|\mathbf{p}_{g-1}\|^{2}$. Hence,
$\mathbf{x}_{g-1}-\frac{\left\langle\mathbf{p}_{g-1},\mathbf{x}_{g-1}\right\rangle}{\|\mathbf{p}_{g-1}\|^{2}}\mathbf{p}_{g-1}$ $=\mathbf{A}^{\dagger}\mathbf{A}\mathbf{x}_{g-1}=\mathbf{x}^{+},$
which leads us to Theorems 4 and 4.
Algorithm 4 Conjugate Residual with Pseudo-inverse Solution
1:a symmetric matrix $\mathbf{A}$ and a vector $\mathbf{b}$
2:Let $k=1$, $\mathbf{r}_{0}=\mathbf{p}_{0}=\mathbf{b}$
3:while $\|\mathbf{Ar}_{k-1}\|\not=0$ do $\triangleright$ Termination Condition
4:     $\alpha_{k-1}=\left\langle\mathbf{r}_{k-1},\mathbf{A}\mathbf{r}_{k-1}\right\rangle/\|\mathbf{A}\mathbf{p}_{k-1}\|^{2}$triangleright Classical CR: Steps 3 - 7
5:     $\mathbf{x}_{k}=\mathbf{x}_{k-1}+\alpha_{k-1}\mathbf{p}_{k-1}$
6:     $\mathbf{r}_{k}=\mathbf{r}_{k-1}-\alpha_{k-1}\mathbf{Ap}_{k-1}$
7:     $\beta_{k-1}=\left\langle\mathbf{r}_{k},\mathbf{A}\mathbf{r}_{k}\right\rangle/\left\langle\mathbf{r}_{k-1},\mathbf{A}\mathbf{r}_{k-1}\right\rangle$
8:     $\mathbf{p}_{k}=\mathbf{r}_{k}+\beta_{k-1}\mathbf{p}_{k-1}$
9:     $k=k+1$
10:end while
11:$\mathbf{x}^{\star}=\mathbf{x}_{k-1}$triangleright Normal Solution
12:$\mathbf{x}^{+}=\mathbf{x}_{k-1}-(\left\langle\mathbf{p}_{k-1},\mathbf{x}_{k-1}\right\rangle/\|\mathbf{p}_{k-1}\|^{2})\mathbf{p}_{k-1}$triangleright Pseudo-inverse Solution
13:return $\mathbf{x}^{+}$
Theorem 4 (Pseudo-inverse Solution with Algorithm 4).
Under Assumption 1, in Algorithm 4, we have $\mathbf{x}^{+}=\mathbf{A}^{\dagger}\mathbf{b}$.
Beyond the fact that CR naturally recovers a normal solution, the procedure for obtaining the pseudo-inverse solution in CR is far more straightforward than that for CG as it merely involves the addition of one simple one-dimensional projection at termination.
Under Assumption 1, we simply have
$\|\mathbf{r}_{t}\|^{2}$ $=\|\mathbf{r}_{t-1}-\alpha_{t-1}\mathbf{Ap}_{t-1}\|^{2}=\|\mathbf{r}_{t-1}\|^{2}+\alpha_{t-1}^{2}\|\mathbf{Ap}_{t-1}\|^{2}-2\alpha_{t-1}\left\langle\mathbf{r}_{t-1},\mathbf{Ap}_{t-1}\right\rangle$
$=\|\mathbf{r}_{t-1}\|^{2}+\frac{\left\langle\mathbf{r}_{t-1},\mathbf{Ar}_{t-1}\right\rangle^{2}}{\|\mathbf{Ap}_{t-1}\|^{2}}-2\frac{\left\langle\mathbf{r}_{t-1},\mathbf{Ar}_{t-1}\right\rangle^{2}}{\|\mathbf{Ap}_{t-1}\|^{2}}=\|\mathbf{r}_{t-1}\|^{2}-\frac{\left\langle\mathbf{r}_{t-1},\mathbf{Ar}_{t-1}\right\rangle^{2}}{\|\mathbf{Ap}_{t-1}\|^{2}}$
$<\|\mathbf{r}_{t-1}\|^{2},$
which implies that the norm of the residuals in Algorithm 2 (or Algorithm 4) monotonically decreasing. On the other hand, MINRES also guarantees a monotonic decrease of the residual norms. The equivalence between CR and MINRES for PD systems is well-studied [14]. Next, we show that, in the absence of an unlucky breakdown, this equivalence indeed extends beyond PD settings to all systems.
3.2.2 Equivalence to MINRES
The minimum residual (MINRES) method of Paige and Saunders [34] is a versatile approach for solving any linear systems/least-squares problems involving symmetric matrices. The iterates are constructed in a way that prevents unlucky breakdowns, guaranteeing the algorithm’s convergence to a normal solution. In fact, Facts 1 and 3 hold true for MINRES without the necessity to exclude any unfortunate situations, such as those specified in Assumption 1.
[…]
$\mathbf{y}_{k}\triangleq\underset{\mathbf{y}\in\mathbb{R}^{k}}{\text{argmin}}\frac{1}{2}\left\|\|\mathbf{b}\|\mathbf{e}_{1}-\tilde{\mathbf{T}}_{k}\mathbf{y}\right\|^{2}.$
In essence, the above problem is solved using the QR factorization of $\tilde{\mathbf{T}}_{k}$. However, by exploiting its sparsity structure, we can simplify the computations and obtain an update procedure of the form $\mathbf{x}_{k}=\mathbf{x}_{k-1}+\tau_{k-1}\mathbf{d}_{k-2}$, for some step size $\tau_{k-1}$ and the update direction $\mathbf{d}_{k}$; see the works [34, 28] for more details.
For positive definite matrices, it is well-known that CR and MINRES, mathematically speaking, generate identical iterates [14]. However, beyond PD settings, such an equivalence, at least to our knowledge, has not been thoroughly studied. In these general settings, the $\mathbf{A}$-weighted product .,.$\left\langle.,.\right\rangle_{\mathbf{A}}$ in CR is no longer a valid inner product and the equivalence between CR and MINRES is no longer automatic. In this section, we study the existence of such an equivalence in indefinite and/or inconsistent settings, and show that, under Assumption 1, both CR and MINRES generate the same sequence of iterates for any symmetric systems.
To that end, we first demonstrate that, just like MINRES, subproblems of CR can also be formulated as 5 for any symmetric matrix $\mathbf{A}$. Let $\mathbf{P}_{k}=\begin{bmatrix}\mathbf{p}_{0}&\cdots&\mathbf{p}_{k-1}\end{bmatrix}$ and $\mathbf{R}_{k}=\begin{bmatrix}\mathbf{r}_{0}&\cdots&\mathbf{r}_{k-1}\end{bmatrix}$ be the matrices whose columns are, respectively, the update directions and the residual vectors from CR. Using Property 1-(a) and noting that $\mathbf{r}_{0}=\mathbf{b}$, we can write $\mathbf{x}=\mathbf{P}_{k}\mathbf{z}$ for $\mathbf{z}\in\mathbb{R}^{k}$ and
$\min_{\mathbf{x}\in\mathcal{K}_{k}(\mathbf{A},\mathbf{b})}\frac{1}{2}\|\mathbf{b}-\mathbf{Ax}\|^{2}$ $=\min_{\mathbf{z}\in\mathbb{R}^{k}}\frac{1}{2}\left\|\mathbf{R}_{k}\mathbf{e}_{1}-\mathbf{AP}_{k}\mathbf{z}\right\|^{2},$
which gives
$\mathbf{z}_{k}$ $\triangleq\underset{\mathbf{z}\in\mathbb{R}^{k}}{\text{argmin}}\frac{1}{2}\left\langle\mathbf{z},\mathbf{P}_{k}^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}\mathbf{A}^{2}\mathbf{P}_{k}\mathbf{z}\right\rangle-\left\langle\mathbf{z},\mathbf{P}_{k}^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}\mathbf{A}\mathbf{R}_{k}\mathbf{e}_{1}\right\rangle.$ (8)
By Property 1-(c), the matrix $\mathbf{D}_{k}\triangleq\mathbf{P}_{k}^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}\mathbf{A}^{2}\mathbf{P}_{k}$ is diagonal, i.e.,
$\mathbf{D}_{k}=\begin{bmatrix}\|\mathbf{Ap}_{0}\|^{2}&&&\\ &\|\mathbf{Ap}_{1}\|^{2}&&\\ &&\ddots&\\ &&&\|\mathbf{Ap}_{k-1}\|^{2}\end{bmatrix}.$
Note that, by Lemma 3, the matrix $\mathbf{D}_{k}$ is always invertible prior to termination. Also, from Properties 1-(a) and 1-(c), it follows that the matrix $\mathbf{K}_{k}\triangleq\mathbf{P}_{k}^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}\mathbf{A}\mathbf{R}_{k}$ is lower triangular, i.e.,
$\mathbf{K}_{k}=\begin{bmatrix}\left\langle\mathbf{r}_{0},\mathbf{Ap}_{0}\right\rangle&0&\cdots&0\\ \left\langle\mathbf{r}_{0},\mathbf{Ap}_{1}\right\rangle&\left\langle\mathbf{r}_{1},\mathbf{Ap}_{1}\right\rangle&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ \left\langle\mathbf{r}_{0},\mathbf{Ap}_{k-1}\right\rangle&\left\langle\mathbf{r}_{1},\mathbf{Ap}_{k-1}\right\rangle&\cdots&\left\langle\mathbf{r}_{k-1},\mathbf{Ap}_{k-1}\right\rangle\end{bmatrix}.$
Optimality condition for 8 implies $\mathbf{z}_{k}=\mathbf{D}_{k}^{-1}\mathbf{K}_{k}\mathbf{e}_{1}$. The vector $\mathbf{K}_{k}\mathbf{e}_{1}$ can be written as,
$\mathbf{K}_{k}\mathbf{e}_{1}=\begin{bmatrix}\left\langle\mathbf{r}_{0},\mathbf{Ap}_{0}\right\rangle\\ \left\langle\mathbf{r}_{0},\mathbf{Ap}_{1}\right\rangle\\ \vdots\\ \left\langle\mathbf{r}_{0},\mathbf{Ap}_{k-1}\right\rangle\end{bmatrix}=\begin{bmatrix}\left\langle\mathbf{r}_{0},\mathbf{Ar}_{0}\right\rangle\\ \beta_{0}\left\langle\mathbf{r}_{0},\mathbf{Ar}_{0}\right\rangle\\ \vdots\\ \prod_{i=1}^{k}\beta_{i-1}\left\langle\mathbf{r}_{0},\mathbf{Ar}_{0}\right\rangle\end{bmatrix}=\begin{bmatrix}\left\langle\mathbf{r}_{0},\mathbf{Ar}_{0}\right\rangle\\ \left\langle\mathbf{r}_{1},\mathbf{Ar}_{1}\right\rangle\\ \vdots\\ \left\langle\mathbf{r}_{k-1},\mathbf{Ar}_{k-1}\right\rangle\end{bmatrix}$
where the last equality follows from the definition of $\beta$ in Algorithm 2. Together, we get
$\mathbf{z}_{k}=\mathbf{D}_{k}^{-1}\mathbf{K}_{k}\mathbf{e}_{1}=\begin{bmatrix}\left\langle\mathbf{r}_{0},\mathbf{Ar}_{0}\right\rangle/\|\mathbf{Ap}_{0}\|^{2}\\ \left\langle\mathbf{r}_{1},\mathbf{Ar}_{1}\right\rangle/\|\mathbf{Ap}_{1}\|^{2}\\ \vdots\\ \left\langle\mathbf{r}_{k-1},\mathbf{Ar}_{k-1}\right\rangle/\|\mathbf{Ap}_{k-1}\|^{2}\end{bmatrix}=\begin{bmatrix}\alpha_{0}\\ \alpha_{1}\\ \vdots\\ \alpha_{k-1}\end{bmatrix},$
[…]
A direct implication of the above result is that CR and MINRES will, in exact arithmetic, generate the same sequence of iterates, $\{\mathbf{x}_{k-1}\}_{k=1}^{k=g}$, for consistent but potentially singular systems.
Lemma 5.
Let $\mathbf{b}\in\text{Range}(\mathbf{A})$. Then, under Assumption 1, CR and MINRES generate the same sequence of iterates.
Proof.
As showed above, both iterates of CR and MINRES correspond to solutions of 5. It suffices to show that, for consistent systems, 5 admits a unique solution for any $1\leq k\leq g$. Let $\mathbf{W}_{k}$ be any basis of $\mathcal{K}_{k}(\mathbf{A},\mathbf{b})$. Defining $\mathbf{x}=\mathbf{W}_{k}\mathbf{y}$ for $\mathbf{y}\in\mathbb{R}^{d}$, we can write 5 as
$\underset{\mathbf{y}\in\mathbb{R}^{k}}{\text{argmin}}\frac{1}{2}\left\|\mathbf{b}-\mathbf{AW}_{k}\mathbf{y}\right\|^{2}=\underset{\mathbf{y}\in\mathbb{R}^{k}}{\text{argmin}}\frac{1}{2}\left\langle\mathbf{y},\mathbf{W}_{k}^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}\mathbf{A}^{2}\mathbf{W}_{k}\mathbf{y}\right\rangle-\left\langle\mathbf{y},\mathbf{W}_{k}^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}\mathbf{Ab}\right\rangle.$
By the assumption on $\mathbf{b}$, $\text{Span}(\mathbf{W}_{k})\subseteq\text{Range}(\mathbf{A})$, which implies $\mathbf{W}_{k}^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}\mathbf{A}^{2}\mathbf{W}_{k}\succ\mathbf{0}$. Hence, the above problem admits a unique solution $\mathbf{y}_{k}$, resulting in a unique $\mathbf{x}_{k}$. ∎
Hence, when $\mathbf{b}\in\text{Range}(\mathbf{A})$, there can be only one unique minimizer of 5 in $\mathcal{K}_{k}(\mathbf{A},\mathbf{b})$ for $1\leq k\leq g$. However, the case for inconsistent systems may appear less straightforward. Specifically, when $\mathbf{b}\not\in\text{Range}(\mathbf{A})$, there may be more than one solution to 5. The crucial step, now, is to show that, despite possibly many solutions to 5, the minimizer from CR is indeed the same as the one from MINRES, i.e., the sequence of iterates $\{\mathbf{x}_{k-1}\}_{k=1}^{k=g}$ for both algorithms, in exact arithmetic, coincide. To do this, we first establish a counterpart of Lemma 4 for MINRES (Lemma 6) and then leverage Lemmas 5, 4 and 6 to establish our claim (Theorem 5). A diagram depicting our proof strategy is given in Figure 1.
[…]
Figure 1: Our proof strategy for establishing equivalence between MINRES and CR. To prove Theorem 5, we use Lemmas 5, 4 and 6.
Lemma 6.
Consider the iterates generated by the procedures MINRES$(\mathbf{A},\tilde{\mathbf{b}})$ and MINRES$(\mathbf{A},\mathbf{b})$, and let where $g$ be the grade of $\mathbf{b}$ with respect to $\mathbf{A}$. For $1\leq k\leq g$, we have $\tilde{\mathbf{x}}_{k-1}=\mathbf{A}^{\dagger}\mathbf{A}\mathbf{x}_{k-1}$.
Proof.
[…]
$\min_{\mathbf{x}\in\mathcal{K}_{k}(\mathbf{A},\mathbf{b})}\|\tilde{\mathbf{b}}-\mathbf{Ax}\|^{2},$
which using $\mathbf{A}=\mathbf{A}\mathbf{A}^{\dagger}\mathbf{A}$ gives
$\min_{\mathbf{x}\in\mathcal{K}_{k}(\mathbf{A},\mathbf{b})}\|\tilde{\mathbf{b}}-\mathbf{A}\mathbf{A}^{\dagger}\mathbf{A}\mathbf{x}\|^{2}.$
Denoting $\tilde{\mathbf{x}}=\mathbf{A}^{\dagger}\mathbf{A}\mathbf{x}$, and using the fact that $\mathbf{A}\mathbf{b}=\mathbf{A}\tilde{\mathbf{b}}$, it follows that $\tilde{\mathbf{x}}\in\mathbf{A}^{\dagger}\mathbf{A}\mathcal{K}_{k}(\mathbf{A},\mathbf{b})=\mathcal{K}_{k}(\mathbf{A},\tilde{\mathbf{b}})$. Hence, 5 can equivalently be written as
$\min_{\tilde{\mathbf{x}}\in\mathcal{K}_{k}(\mathbf{A},\tilde{\mathbf{b}})}\|\tilde{\mathbf{b}}-\mathbf{A}\tilde{\mathbf{x}}\|^{2},$
[…]
Theorem 5.
Under Assumption 1, CR and MINRES generate the same sequence of iterates.
Proof.
[…]
$\mathbf{x}_{k-1}^{\text{MINRES}}$ $=\sum_{i=1}^{k-1}a_{i-1}\mathbf{A}^{i-1}\mathbf{b},\quad\text{and}\quad\mathbf{x}_{k-1}^{\text{CR}}=\sum_{i=1}^{k-1}b_{i-1}\mathbf{A}^{i-1}\mathbf{b}.$
Using Lemmas 5, 4 and 6, we have $\mathbf{A}^{\dagger}\mathbf{A}\mathbf{x}_{k-1}^{\text{MINRES}}=\mathbf{A}^{\dagger}\mathbf{A}\mathbf{x}_{k-1}^{\text{CR}}$, which implies
$\sum_{i=1}^{k-1}(a_{i-1}-b_{i-1})\mathbf{A}^{i-1}\tilde{\mathbf{b}}=\mathbf{0}.$
Now, from linear independence of $\{\mathbf{A}^{i-1}\tilde{\mathbf{b}}\}_{i=1}^{k-1}$, we get $a_{i-1}=b_{i-1}$ for $11\leq i\leq k-$ 1. ∎
Theorem 5 implies that all properties established for MINRES in prior works, such as non-positive curvature detection, monotonicity, and lifting procedures in [28, 26], are also readily applicable to CR.
After establishing the equivalence between MINRES and CR under Assumption 1, it is of theoretical interest to investigate the behavior of MINRES when CR experiences an unlucky breakdown. To address this, we present the following two lemmas that characterize this behavior.
Lemma 7.
In MINRES/CR, for $k\leq g-1$, the residual vectors $\{\mathbf{r}_{0},\mathbf{r}_{1},\cdots,\mathbf{r}_{k}\}$ are linearly independent if and only if $\langle\mathbf{r}_{i-1},\mathbf{Ar}_{i-1}\rangle\neq 0$ for $1\leq i\leq k$.
Proof.
[…]
Note that if $\left\langle\mathbf{r}_{k},\mathbf{Ar}_{k}\right\rangle=0$ for the first time, the residual vectors $\{\mathbf{r}_{0},\mathbf{r}_{1},\cdots,\mathbf{r}_{k}\}$ are still linearly independent.
Lemma 8.
In MINRES/CR, let $k\leq g-1$. The condition $\langle\mathbf{r}_{k-1},\mathbf{A}\mathbf{r}_{k-1}\rangle=0$ (indicating an unlucky breakdown for CR) occurs for the first time if and only if the Lanczos tridiagonal matrix $\mathbf{T}_{k}$ becomes singular for the first time.
Proof.
Let $\mathbf{V}_{k}\in\mathbb{R}^{d\times k}$ be the Lanczos matrix with orthonormal columns, generated within MINRES, that span $\mathcal{K}_{k}(\mathbf{A},\mathbf{b})$. Recall that $\mathbf{V}_{k}^{\top}\mathbf{AV}_{k}=\mathbf{T}_{k}\in\mathbb{R}^{k\times k}$, where $\mathbf{T}_{k}$ is tridiagonal. Let $\mathbf{R}_{k}=\begin{bmatrix}\mathbf{r}_{0}&\mathbf{r}_{1}&\cdots&\mathbf{r}_{k-1}\end{bmatrix}=\mathbf{V}_{k}\mathbf{Z}_{k}$ for some $\mathbf{Z}_{k}\in\mathbb{R}^{k\times k}$. Consider the following expression,
$\begin{bmatrix}\left\langle\mathbf{r}_{0},\mathbf{Ar}_{0}\right\rangle&&&\\ &\left\langle\mathbf{r}_{1},\mathbf{Ar}_{1}\right\rangle&&\\ &&\ddots&\\ &&&\left\langle\mathbf{r}_{k-1},\mathbf{Ar}_{k-1}\right\rangle\end{bmatrix}=\mathbf{R}_{k}\mathbf{AR}_{k}=\mathbf{Z}_{k}^{\top}\mathbf{V}_{k}^{\top}\mathbf{AV}_{k}\mathbf{Z}_{k}=\mathbf{Z}_{k}^{\top}\mathbf{T}_{k}\mathbf{Z}_{k}.$
where the first equality follows from Property 1-(c). By Lemma 7, as long as $\left\langle\mathbf{r}_{i},\mathbf{Ar}_{i}\right\rangle\neq 0$ for $20\leq i\leq k-$ 2, the columns of $\mathbf{R}_{k}$ are linearly independent and $\mathbf{Z}_{k}$ must be full rank. Hence, $\left\langle\mathbf{r}_{k-1},\mathbf{Ar}_{k-1}\right\rangle=0$ if and only if $\mathbf{T}_{k}$ is singular. ∎
4 Numerical Experiments
In this section, we conduct a series of experiments to verify aspects of the theory developed in this paper and to explore the empirical performance of the methods discussed in various settings. While numerical inaccuracies due to finite precision arithmetic in iterative methods have long been well known and thoroughly studied, our main focus is to highlight the additional and unexpectedly catastrophic instability of CG in inconsistent settings–behavior not typically observed in consistent settings–and compare it to the CR alternative. In particular, we set out to demonstrate the following:
  1. 1.
    CR is generally more stable than CG in most cases, regardless of consistency.
  2. 2.
    Unlike CR, CG exhibits extreme instability under inconsistent systems.
  3. 3.
    Unlike CR, CG can be unstable even for PD matrices with a few small eigenvalues.
In Section 4.1, using synthetic examples, our aim is to empirically verify Lemmas 1 and 1. Additionally, we will investigate the stability of these methods and identify instances where the theoretical predictions deviate significantly from practical observations. Moving to Section 4.2, our focus remains on the same objectives, but within the context of real-world applications. It is crucial to note that our primary goal here is to observe the behavior of these methods in their original form, without incorporating any additional enhancements such as reorthogonalization strategies (if applicable).
Algorithms are terminated either upon reaching a maximum number of iterations or when a specific termination condition is met. For CG, this condition is chosen as $\min\{\|\mathbf{r}_{k}\|/\|\mathbf{b}\|,\|\mathbf{Ap}_{k}\|/\|\mathbf{Ab}\|\}$ (cf. Facts 1 and 1), and for CR/MINRES, it is $\min\{\|\mathbf{r}_{k}\|/\|\mathbf{b}\|,\|\mathbf{Ar}_{k}\|/\|\mathbf{Ab}\|\}$ (cf. Facts 1 and 3).
4.1 Synthetic Problems
In this section, we experiment with a series of simple and synthetic diagonal matrices to verify Lemmas 1 and 1 and to explore the stability of these methods in several instances. To validate Lemmas 1 and 1, we focus on small-size problems in Section 4.1.1 to avoid any unwanted numerical issues. Subsequently, we extend our analysis to larger-size matrices in Sections 4.1.3 and 4.1.2 to investigate the stability of these methods across a substantial number of iterations.
For our experiments, consider various matrices that are generated according to the following template:
  • PSD Matrix: A diagonal matrix generated by sampling from a uniform distribution, with $m$ zero diagonals.
  • Indefinite Matrix: A diagonal matrix generated by sampling from a normal distribution, with $m$ zero diagonals.
  • (Slightly) Indefinite Matrix: A diagonal matrix sampled from a uniform distribution, with $m$ zero diagonals. Additionally, $\lceil 0.1\times g\rceil$ of these diagonals have their signs flipped to negative, where $g$ represents the grade.
The vector $\mathbf{b}\in\mathbb{R}^{d}$ is sampled from a normal distribution. The examples in this section are constructed in such a way that they are always inconsistent, unless otherwise stated.
4.1.1 CG at Termination: Validating Lemmas 1 and 1
In this section, we validate Lemmas 1 and 1. For this, we consider $d=10$, and generate six matrices by randomly zeroing out $m=5,4,3,2,1$ and 00 diagonal elements of $\mathbf{A}$. The results are depicted in Figures 2 and 3 where we consider PSD and indefinite settings, respectively. In all these examples, as predicted by Lemma 1, $\|\mathbf{Ap}_{k}\|$ indeed vanishes at termination, while $\|\mathbf{Ar}_{k}\|$ remains bounded away from zero in all but consistent case, validating Theorem 1.
Figure 2: Experiments for Section 4.1.1 to verify Lemmas 1 and 1. The matrices are PSD with $d=10$. From the top left to the bottom right, the grades are 6, 7, 8, 9, 10 and 10. The bottom right is a positive definite matrix.
Figure 3: Experiments for Section 4.1.1 to verify Lemmas 1 and 1. The matrices are indefinite with $d=10$. From the top left to the bottom right, the grades are 6, 7, 8, 9, 10 and 10. The bottom right is a full rank indefinite matrix.
4.1.2 Numerical Instability of CG
In this section, we consider larger size matrices than those in Section 4.1.1 to explore the behavior of CG across a larger number of iterations. Specifically, we set $d=100$ and generate a series of matrices by zeroing out $m=90,80,60,20,1$ and 00 diagonal entries. The maximum number of iterations allowed is capped at $g$, i.e., the corresponding grade for each example, and the tolerance for the termination condition $\min\{\|\mathbf{r}_{k}\|/\|\mathbf{b}\|,\|\mathbf{Ap}_{k}\|/\|\mathbf{Ab}\|\}$ is $10^{-8}$. The results are gathered in Figures 4, 5 and 6, corresponding to PSD, indefinite, and slightly indefinite settings, respectively. In all these examples, when $g$ is small, the theoretical prediction of Lemma 1 aligns with the observed behavior. However, as $g$ increases, necessitating more iterations from CG, a clear discrepancy emerges. Unless the system is consistent, the implications outlined in Lemma 1 do not accurately reflect practical outcomes, highlighting the presence of substantial numerical inaccuracies within CG iterations. Nonetheless, an interesting observation from Figures 6 and 5 is that the presence of negative eigenvalues appears to induce a form of oscillatory behavior in CG. While the depicted quantities remain bounded away from zero, they do not exhibit the pronounced growth observed in Figure 4.
Figure 4: Experiments for Section 4.1.2 to explore instabilities within CG. The matrices are PSD with $d=100$. From top left to bottom right, the grades are 11, 21, 41, 81, 100 and 100. The bottom right is a positive definite matrix for which CG terminates at iteration 59 with $\|\mathbf{r}_{k}\|/\|\mathbf{b}\|<10^{-8}$.
Figure 5: Experiments for Section 4.1.2 to explore instabilities within CG. The matrices are indefinite with $d=100$. From top left to bottom right, the grades are 11, 21, 41, 81, 100 and 100. The bottom right is an indefinite matrix with full rank for which CG reaches the maximum allowable iteration.
Figure 6: Experiments for Section 4.1.2 to explore instabilities within CG. The matrices are slightly indefinite with $d=100$. From the top left to the bottom right, the number of negative diagonal entries is $1,2,4,8,10$ and $10$, while the grades are as in Figure 5. The bottom right is a slightly indefinite matrix with full rank for which CG is terminated after the termination condition is triggered.
4.1.3 Numerical (In)Stability of CR/MINRES
In light of the observations in Section 4.1.2 concerning the apparent numerical instabilities of CG, our focus now shifts to examining the stability, or lack thereof, of CR and MINRES. In the absence of appropriate preconditioning or stabilization strategies like reorthogonalization, iterative procedures, such as CG, CR, or MINRES, inevitably encounter numerical challenges when applied to large and ill-conditioned systems. Nevertheless, in this section, we aim to explore and identify which of these methods, if any, demonstrate a visibly more adverse and unstable behavior.
For the experiments in this section, we set $d=1,000$ and $m=800$. The maximum number of iterations is capped at $800$, and the termination condition is triggered at the threshold $10^{-8}$. To assess the extent of numerical inaccuracies, we monitor the fundamental theoretical properties of these methods, specifically orthogonality and conjugacy of the relevant vectors, as per Property 1-(b) and Property 1-(c). For CG, the deviation from the properties outlined in Property 1-(b) is calculated at each iteration. This involves gathering the update directions $\mathbf{p}_{k-1}$ and the normalized residual vectors $\mathbf{r}_{k-1}/\left\|\mathbf{r}_{k-1}\right\|$ as columns of matrices $\mathbf{P}_{k}$ and $\mathbf{R}_{k}$, respectively. The spectral norm of the matrices $\mathbf{I}_{k}-\mathbf{R}_{k}^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}\mathbf{R}_{k}$ and $\text{diag}(\{\left\langle\mathbf{p}_{i},\mathbf{Ap}_{i}\right\rangle\}_{i=0}^{i=k})-\mathbf{P}_{k}^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}\mathbf{AP}_{k}$ is then calculated, where “diag” refers to a diagonal matrix constructed from the given set of values. Following the properties outlined in Property 1-(c), similar quantities are also calculated for CR and MINRES, namely $\mathbf{I}_{k}-(\mathbf{AP}_{k})^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}\mathbf{AP}_{k}$ (here the vectors $\mathbf{Ap}_{k-1}$ are normalized) and $\text{diag}(\{\left\langle\mathbf{r}_{i},\mathbf{Ar}_{i}\right\rangle\}_{i=0}^{i=k})-\mathbf{R}_{k}^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}\mathbf{AR}_{k}$. In the absence of numerical issues, these quantities should all be evaluated to zero.
In addition, we also monitor $\|(\mathbf{b}-\mathbf{Ax}_{k})-\mathbf{r}_{k}\|$ where $\mathbf{r}_{k}$ is the residual vector calculated naturally as part of the algorithm, for example, Line 5 in Algorithms 1 and 2 (a similar quantity also arise as part of MINRES; see [28]). Theoretically, the residual calculated within the algorithm must align with that calculated from the iterate as $\mathbf{b}-\mathbf{Ax}_{k}$. In practice, however, since they are obtained through different means, we expect some discrepancies between them due to numerical errors. Nevertheless, methods that are inherently more stable should exhibit less severe discrepancies.
The results of the experiments for this section are gathered Figures 7, 8 and 9. In stark contrast to CR and MINRES, CG exhibits a significant degree of numerical instability in Figure 7, reaching a level that can be categorized as catastrophic. These observations align well with the findings in [23] who suggests that CG under inconsistent systems could lead to divergence, though it does not provide any more insights. Furthermore, in Figures 8 and 9, although all methods experience some loss of orthogonality and conjugacy in their underlying vectors, CR and MINRES demonstrate slightly more stable numerical behavior. This is evident in the quantity $\|(\mathbf{b}-\mathbf{Ax}_{k})-\mathbf{r}_{k}\|$, and they make relatively satisfactory progress toward obtaining a normal solution with a relatively small residual from the normal equations.
CG
CR
MINRES
Figure 7: Experiments for Section 4.1.3 to compare numerical (in)stability among CG, CR, and MINRES. The matrix is PSD with $d=1000$ and $m=800$. Surprisingly, the relative residual and $\|(\mathbf{b}-\mathbf{Ax}_{k})-\mathbf{r}_{k}\|$ for CG reach catastrophically large levels. In contrast, not only do both CR and MINRES perform almost identically as predicted, but they also exhibit significantly more stability compared to CG.
CG
CR
MINRES
Figure 8: Experiments for Section 4.1.3 to compare numerical (in)stability among CG, CR, and MINRES. The matrix is indefinite with $d=1000$ and $m=800$. While all methods exhibit some loss of orthogonality and conjugacy in the relevant vectors, CR and MINRES show slightly more stability as measured by $\|(\mathbf{b}-\mathbf{Ax}_{k})-\mathbf{r}_{k}\|$. As in Figures 6 and 5, the presence of negative eigenvalues appears to induce a form of oscillatory behavior in CG and the depicted quantities do not exhibit the pronounced growth observed in Figure 7.
CG
CR
MINRES
Figure 9: Experiments for Section 4.1.3 to compare numerical (in)stability among CG, CR, and MINRES. The matrix is slightly indefinite with $d=1000$, $m=800$, and only $80$ negative diagonal entries. While all methods exhibit some loss of orthogonality and conjugacy in the relevant vectors, CR and MINRES show slightly more stability as measured by $\|(\mathbf{b}-\mathbf{Ax}_{k})-\mathbf{r}_{k}\|$. In fact, they both converge to a solution with a relatively small residual of the normal equation. Again, the presence of negative eigenvalues appears to induce a form of oscillatory behavior in CG. Yet, compared with Figure 8 where the matrix contained more negative eigenvalues, the depicted quantities for CG exhibit a more pronounced growth.
4.2 Real-world Applications
In this section, we now evaluate the behavior of CG, CR, and MINRES in some real-world applications. Namely, we consider problems from the numerical solution of partial differential equations (Section 4.2.1), computer vision (Section 4.2.2), and statistical modeling (Section 4.2.3).
4.2.1 Poisson Equation with Neumann Boundary Condition
Consider numerically solving a 2D Poisson equation with pure Neumann boundary conditions,
$-\Delta u$ $=f(x,y)\quad\text{where}\quad(x,y)\in\Omega$
$\left\langle\nabla u,\mathbf{n}\right\rangle$ $=g(x,y)\quad\text{where}\quad(x,y)\in\partial\Omega,$
where $u:\Omega\subseteq\mathbb{R}^{2}\to\mathbb{R}$ is the unknown function. Here, $\Delta u$, $\nabla u$ and $\mathbf{n}$ are the Laplace operator of $u$, the gradient of $u$, and the normal vector to the boundary $\partial\Omega$ (the exterior surface), respectively. We use a five-point stencil on uniform staggered grids to formulate the finite difference of the Laplace operator of $u$. The resulting Laplacian matrix of this particular problem is known to be PSD and rank deficient [44, 37].
In our experiments, we generate $f$ and $g$ by letting $u(x,y)=\sin(\sqrt{x^{2}+y^{2}})$ as the true, and considered unknown, underlying function; see Figure 10. The domain is centered at $(0.001,0.001)$ with size $(x,y)\in\Omega\triangleq[-10+0.001,10+0.001]\times[-10+0.001,10+0.001]$. The slight off-center positioning is attributed to the undefined derivatives of $u$ at $(0,0)$. We partition the domain into grids of size $512\times 512$, resulting in a sparse Laplace matrix with dimensions $(512+1)^{2}\times(512+1)^{2}$. The maximum number of iterations and the termination condition are set to $2,000$ and $\|\mathbf{Ar}_{k}\|/\|\mathbf{Ab}\|\leq 10^{-10}$, respectively.
The results are depicted in Figures 11 and 12. In Figure 11, unlike CR and MINRES, CG fails to attain the threshold required for the termination condition, prior to reaching the maximum number of iterations. Not only, do both CR and MINRES perform identically, but also they obtain a solution with significantly superior quality compared to that obtained from CG upon termination. Figure 12 illustrates the quality of the recovered solutions from all methods at iteration 359, the point where CG achieves its lowest residual. While the residuals for CR and MINRES are lower than those for CG at this stage, the recovered solution for them exhibits a larger error.
Figure 10: The true, yet unknown, function $u(x,y)=\sin(\sqrt{x^{2}+y^{2}})$ for generating the experiments of Section 4.2.1.
CG
CR
MINRES
Figure 11: Experiments for Section 4.2.1. Both CR and MINRES behave identically and terminate at iteration 1640 with $\|\mathbf{r}_{k}\|/\|\mathbf{b}\|\approx 1.12\times 10^{-3}$, $\|\mathbf{Ar}_{k}\|/\|\mathbf{Ab}\|\approx 9.91\times 10^{-11}$ and $\|\mathbf{x}_{\text{true}}-\mathbf{x}_{k}\|/\|\mathbf{x}_{\text{true}}\|\approx 0.0841$. CG never attains the termination condition threshold prior to reaching the maximum number of iterations. At termination, CG obtains $\|\mathbf{r}_{k}\|/\|\mathbf{b}\|\approx 5.06\times 10^{5}$, $\|\mathbf{Ar}_{k}\|/\|\mathbf{Ab}\|\approx 6.43\times 10^{5}$, $\|\mathbf{Ap}_{k}\|/\|\mathbf{Ab}\|\approx 4.9\times 10^{5}$ and $\|\mathbf{x}_{\text{true}}-\mathbf{x}_{k}\|/\|\mathbf{x}_{\text{true}}\|\approx 8.7\times 10^{13}$.
Figure 12: Experiments for Section 4.2.1. CG attains its minimum relative residual at iteration 359 and the recovered solutions for all three methods at this specific iteration are depicted. At this iteration, CG obtains $\|\mathbf{r}_{k}\|/\|\mathbf{b}\|\approx 1.22\times 10^{-2}$, $\|\mathbf{Ar}_{k}\|/\|\mathbf{Ab}\|\approx 1.98\times 10^{-2}$, $\|\mathbf{Ap}_{k}\|/\|\mathbf{Ab}\|\approx 1.15\times 10^{-2}$ and $\|\mathbf{x}_{\text{true}}-\mathbf{x}_{k}\|/\|\mathbf{x}_{\text{true}}\|\approx 0.0992$, while for CR and MINRES we observe $\|\mathbf{r}_{k}\|/\|\mathbf{b}\|\approx 2.25\times 10^{-3}$, $\|\mathbf{Ar}_{k}\|/\|\mathbf{Ab}\|\approx 3.92\times 10^{-4}$ and $\|\mathbf{x}_{\text{true}}-\mathbf{x}_{k}\|/\|\mathbf{x}_{\text{true}}\|\approx 0.1212$.
4.2.2 Gaussian Smoothing
Next, we consider the behavior of CG, CR, and MINRES in the context of an image deblurring problem [25]. For a given image $\mathbf{x}$, we apply a linear Gaussian smoothing / blurring operator $\mathbf{A}$, as a 2D convolutional operator, on the image and add some noise $\mathbf{e}$ sampled from a uniform distribution, i.e., $\mathbf{b}=\mathbf{A}\mathbf{x}+\mathbf{e}$. We follow a similar experimental setup as the numerical experiment in the work[26]. The true image and blurry version are given in Figure 13. We remind that though $\mathbf{A}$ typically has full rank, it often contains several small singular values, which make it almost numerically singular. We do not use any particular regularization and simply apply CG, CR and MINRES to attempt to recover the original image. The maximum number of iterations and the termination condition are set to $300$ and $\|\mathbf{Ar}_{k}\|/\|\mathbf{Ab}\|\leq 10^{-5}$, respectively.
The results are presented in Figures 14 and 15. We stress that even when the matrix $\mathbf{A}$ is PD with a few small eigenvalues, it can be nearly numerically singular. Consequently, CG can exhibit behavior similar to that observed in inconsistent systems and may fail to achieve the required termination threshold before reaching the maximum number of iterations. In contrast, both CR and MINRES, as evident from Figures 14 and 15, demonstrate greater stability and terminate once $\|\mathbf{Ar}_{k}\|/\|\mathbf{Ab}\|\leq 10^{-5}$. Clearly, the image recovered by CG at termination is of significantly lower quality compared to those obtained by CR and MINRES. Similar to the previous section, Figure 15 showcases the best performance of CG — the image recovered by CG at the iteration with the smallest relative residual — compared with the images from CR and MINRES at the same iteration. At this particular iteration, while CG produces a slightly improved image, it remains visibly blurry with a larger relative error than those obtained by CR/MINRES at termination.
Figure 13: The true and the noisy blurred images used for the experiments of Section 4.2.2.
CG
CR
MR
Figure 14: Experiments for Section 4.2.2. The quantities plotted correspond to the red channel of the images. CG terminates at iteration 300 with $\|\mathbf{r}_{k}\|/\|\mathbf{b}\|\approx 0.511$, $\|\mathbf{Ar}_{k}\|/\|\mathbf{Ab}\|\approx 0.323$, $\|\mathbf{Ap}_{k}\|/\|\mathbf{Ab}\|\approx 0.184$ and $\|\mathbf{x}_{\text{true}}-\mathbf{x}_{k}\|/\|\mathbf{x}_{\text{true}}\|\approx 1.8\times 10^{5}$. CR and MINRES both satisfy the termination condition at iteration 30 with $\|\mathbf{r}_{k}\|/\|\mathbf{b}\|\approx 1.89\times 10^{-3}$, $\|\mathbf{Ar}_{k}\|/\|\mathbf{Ab}\|\approx 9.76\times 10^{-6}$ and $\|\mathbf{x}_{\text{true}}-\mathbf{x}_{k}\|/\|\mathbf{x}_{\text{true}}\|\approx 0.2376$.
Figure 15: Experiments for Section 4.2.2. CG attains its minimum relative residual at iteration 5 and the recovered solutions for all three methods at this specific iteration are depicted. At this iteration, CG obtains $\|\mathbf{r}_{k}\|/\|\mathbf{b}\|\approx 3.9\times 10^{-3}$, $\|\mathbf{Ar}_{k}\|/\|\mathbf{Ab}\|\approx 2.33\times 10^{-3}$, $\|\mathbf{Ap}_{k}\|/\|\mathbf{Ab}\|\approx 1.33\times 10^{-3}$ and $\|\mathbf{x}_{\text{true}}-\mathbf{x}_{k}\|/\|\mathbf{x}_{\text{true}}\|\approx 0.249$, while for CR and MINRES, we observe $\|\mathbf{r}_{k}\|/\|\mathbf{b}\|\approx 2.55\times 10^{-3}$, $\|\mathbf{Ar}_{k}\|/\|\mathbf{Ab}\|\approx 5.66\times 10^{-4}$ and $\|\mathbf{x}_{\text{true}}-\mathbf{x}_{k}\|/\|\mathbf{x}_{\text{true}}\|\approx 0.2562$.
4.2.3 Kernel Regression
Finally, we consider kernel regression [17], which appears frequently in machine learning applications. Kernel regression is an example of non-parametric regression where given training data $\mathcal{D}=\{\mathbf{a}_{i},b_{i}\}_{i=1}^{n}\subset\mathbb{R}^{p}\times\mathbb{R}$, we aim to find a prediction function $f:\mathbb{R}^{p}\to$ that is a solution to the following minimization problem
$\min_{f\in\mathcal{H}}\left(f(\mathbf{a}_{i})-b_{i}\right)^{2},$
where $\mathcal{H}$ is some appropriate reproducing kernel Hilbert space (RKHS) with the kernel function $\kappa:^{p}\times^{p}\to$ [32]. It can be shown that any solution to the above problem can be written as $f=g+h$ where $g=\sum_{j=1}^{n}x_{j}\kappa(\mathbf{a}_{j},.)$ and $h\perp\text{Span}\{\kappa(\mathbf{a}_{j},.)\}_{j=1}^{n}$, leading to the least-squares regression problem
$\min_{\mathbf{x}\in\mathbb{R}^{n}}\left\|\mathbf{b}-\mathbf{A}\mathbf{x}\right\|^{2},$
where $\mathbf{A}\in\mathbb{R}^{n\times n}$ is the positive (semi-)definite kernel matrix such that $\mathbf{A}_{i,j}=\kappa(\mathbf{a}_{i},\mathbf{a}_{j})$ and $\mathbf{b}\in^{n}$ with $\mathbf{b}_{i}=b_{i}$.
We use Wine Quality dataset, comprising $6,497$ training examples with $p=10$ [5]. The dataset is randomly split into $80\%$ training (thus, $n=5,197$) and $20\%$ validation sets. Employing the radial basis function kernel $\kappa(\mathbf{a},\mathbf{a}^{\prime})=\exp{(-L\|\mathbf{a}-\mathbf{a}^{\prime}\|_{2}^{2})}$ with $L=10^{-4}$ results in the positive semi-definite kernel matrix $\mathbf{A}$ of rank $438$. The maximum number of iterations and the termination condition are set to $2,000$ and $\|\mathbf{Ar}_{k}\|/\|\mathbf{Ab}\|\leq 10^{-7}$, respectively. To assess the predictive performance of the recovered model, we measure the mean squared error (MSE) for the test set at each iteration. The results, depicted in Figure 16, clearly show that both CR and MINRES greatly outperform CG, not only in metrics related to the training set but also in terms of predictive performance as measured by the MSE on the test set.
Figure 16: Experiments for Section 4.2.3. CR and MINRES both attain $\|\mathbf{r}_{k}\|/\|\mathbf{b}\|\approx 0.0835$, $\|\mathbf{Ar}_{k}\|/\|\mathbf{Ab}\|\approx 7.27\times 10^{-8}$ and the validation MSE of 0.025 at 211 and 315 iterations. Whereas, for CG, $\|\mathbf{r}_{k}\|/\|\mathbf{b}\|\approx 3.81$, $\|\mathbf{Ar}_{k}\|/\|\mathbf{Ab}\|\approx 0.976$, $\|\mathbf{Ap}_{k}\|/\|\mathbf{Ab}\|\approx 0.596$ and the validation MSE is 1.1 at iteration 2000.
5 Conclusion
We theoretically and empirically studied the behavior of CG and CR, as two prominent examples of the conjugate directions methods, for solving symmetric but inconsistent systems. We established a fundamental limitation in classical CG in terms of its inability to recover any solution to the normal equation. To remedy this, we proposed a modification that ensures convergence to the pseudo-inverse solution. We then showed that a minor modification to CR at the very last iteration ensures convergence to the pseudo-inverse solution. We further established that CR and MINRES are both essentially equivalent in that they, mathematically speaking, generate the same sequence of iterates. Lastly, we explored the empirical performance of CG, CR, and MINRES in a variety of settings. Contrary to conventional wisdom and in stark contrast to CR/MINRES, we showcased that CG can exhibit numerical instability in a borderline catastrophic manner across scenarios involving inconsistent systems.
Appendix A Unifying Framework
In this section, we present some technical results for Algorithms 1 and 2 that serve as foundational tools in all of our derivations in this paper. To simplify the exposition, we adopt a more general framework depicted in Algorithm 5, which encapsulates both Algorithms 1 and 2 at the same time. Specifically, considering $s=0$ and $s=1$ in Algorithm 5 amount to Algorithm 1 and Algorithm 2, respectively. In essence, it is shown that all common properties of CG and CR, well-known for positive definite settings, continue to hold under indefinite matrices. The proofs are straightforward adaptation of the PD, found in many classical resources, e.g., [33, 41].
Algorithm 5 Conjugate Direction Methods ($s=0$ is CG while $s=1$ is CR)
1:$\mathbf{A}$ and $\mathbf{b}$, and $\mathbf{x}_{0}$
2:Let $k=1$, $\mathbf{r}_{0}=\mathbf{p}_{0}=\mathbf{b}-\mathbf{Ax}_{0}$
3:while Not Terminated do
4:     $\alpha_{k-1}^{(s)}=\left\langle\mathbf{r}_{k-1},\mathbf{A}^{s}\mathbf{r}_{k-1}\right\rangle/\left\langle\mathbf{p}_{k-1},\mathbf{A}^{s+1}\mathbf{p}_{k-1}\right\rangle$
5:     $\mathbf{x}_{k}=\mathbf{x}_{k-1}+\alpha_{k-1}^{(s)}\mathbf{p}_{k-1}$
6:     $\mathbf{r}_{k}=\mathbf{r}_{k-1}-\alpha_{k-1}^{(s)}\mathbf{Ap}_{k-1}$
7:     $\beta_{k-1}^{(s)}=\left\langle\mathbf{r}_{k},\mathbf{A}^{s}\mathbf{r}_{k}\right\rangle/\left\langle\mathbf{r}_{k-1},\mathbf{A}^{s}\mathbf{r}_{k-1}\right\rangle$
8:     $\mathbf{p}_{k}=\mathbf{r}_{k}+\beta_{k-1}^{(s)}\mathbf{p}_{k-1}$
9:     $k=k+1$
10:end while
11:return $\mathbf{x}_{k}$
Lemma 9.
Let $\{\mathbf{r}_{0},\cdots,\mathbf{r}_{k-1}\}$ and $\{\mathbf{p}_{0},\cdots,\mathbf{p}_{k-1}\}$ be the set of vectors generated by Algorithm 5 up to iteration $k-1$, we have
$\text{Span}(\mathbf{p}_{0},\mathbf{p}_{1},\cdots,\mathbf{p}_{k-1})=\text{Span}(\mathbf{r}_{0},\mathbf{r}_{1},\cdots,\mathbf{r}_{k-1})=\text{Span}(\mathbf{r}_{0},\mathbf{Ar}_{0},\cdots,\mathbf{A}^{k-1}\mathbf{r}_{0}).$
Proof.
We prove by induction. For $j=0$, we trivially have $\text{Span}(\mathbf{p}_{0})=\text{Span}(\mathbf{r}_{0})$. Now assume
$\text{Span}(\mathbf{p}_{0},\mathbf{p}_{1},\cdots,\mathbf{p}_{i-1})=\text{Span}(\mathbf{r}_{0},\mathbf{r}_{1},\cdots,\mathbf{r}_{i-1})$ $=\text{Span}(\mathbf{r}_{0},\mathbf{Ar}_{0},\cdots,\mathbf{A}^{i-1}\mathbf{r}_{0})$
for some $i$. We show the claim for $i+1$. Since $\mathbf{p}_{i-1}\in\text{Span}(\mathbf{r}_{0},\mathbf{Ar}_{0},\cdots,\mathbf{A}^{i-1}\mathbf{r}_{0})$, we have $\mathbf{Ap}_{i-1}\in\text{Span}(\mathbf{Ar}_{0},\cdots,\mathbf{A}^{i}\mathbf{r}_{0})$, which from $\mathbf{r}_{i}=\mathbf{r}_{i-1}-\alpha_{i-1}^{(s)}\mathbf{Ap}_{i-1}$ implies $\mathbf{r}_{i}\in\text{Span}(\mathbf{r}_{0},\mathbf{Ar}_{0},\cdots,\mathbf{A}^{i}\mathbf{r}_{0})$. So, $\text{Span}(\mathbf{r}_{0},\mathbf{r}_{1},\cdots,\mathbf{r}_{i})\subset\text{Span}(\mathbf{r}_{0},\cdots,\mathbf{A}^{i}\mathbf{r}_{0})$. However, from $\mathbf{A}^{i}\mathbf{r}_{0}=\mathbf{A}(\mathbf{A}^{i-1}\mathbf{r}_{0})$, we have
$\mathbf{A}^{i}\mathbf{r}_{0}$ $\in\mathbf{A}\text{Span}(\mathbf{p}_{0},\mathbf{p}_{1},\cdots,\mathbf{p}_{i-1})=\text{Span}(\mathbf{Ap}_{0},\mathbf{Ap}_{1},\cdots,\mathbf{Ap}_{i-1})$
$=\text{Span}((\mathbf{r}_{1}-\mathbf{r}_{0}),(\mathbf{r}_{2}-\mathbf{r}_{1}),\cdots,(\mathbf{r}_{i-1}-\mathbf{r}_{i}))=\text{Span}(\mathbf{r}_{0},\mathbf{r}_{1},\cdots,\mathbf{r}_{i}).$
Together, we get $\text{Span}(\mathbf{b},\mathbf{Ab},\cdots,\mathbf{A}^{j}\mathbf{b})=\text{Span}(\mathbf{r}_{1},\mathbf{r}_{2},\cdots,\mathbf{r}_{j+1})$. Finally,
$\text{Span}(\mathbf{p}_{0},\mathbf{p}_{1},\cdots,\mathbf{p}_{i})$ $=\text{Span}(\mathbf{p}_{0},\mathbf{p}_{1},\cdots,\mathbf{p}_{i-1},\mathbf{r}_{i})=\text{Span}(\mathbf{r}_{0},\mathbf{Ar}_{0},\cdots,\mathbf{A}^{i-1}\mathbf{r}_{0},\mathbf{r}_{i})$
$=\text{Span}(\mathbf{r}_{0},\mathbf{r}_{1},\cdots,\mathbf{r}_{i},\mathbf{r}_{i-1}),$
where the first equality is by $\mathbf{p}_{i}=\mathbf{r}_{i}-\beta_{i}^{(s)}\mathbf{p}_{i-1}$
Lemma 10.
Let $\{\mathbf{r}_{0},\cdots,\mathbf{r}_{k-1}\}$ and $\{\mathbf{p}_{0},\cdots,\mathbf{p}_{k-1}\}$ be the set of vectors generated by Algorithm 5 up to iteration $k-1$, for any $10\leq i<j\leq k-$ 1, we have
$\left\langle\mathbf{r}_{i},\mathbf{A}^{s}\mathbf{r}_{j}\right\rangle$ $=0,$ (9a)
$\left\langle\mathbf{p}_{i},\mathbf{A}^{s+1}\mathbf{p}_{j}\right\rangle$ $=0,$ (9b)
Proof.
Without loss of generality, assume $k>1$, we prove the claims by induction. For $i+1=j=1$, $\left\langle\mathbf{r}_{0},\mathbf{A}^{s}\mathbf{r}_{1}\right\rangle=\left\langle\mathbf{r}_{0},\mathbf{A}^{s}(\mathbf{r}_{0}-\alpha_{0}^{(s)}\mathbf{A}\mathbf{p}_{0})\right\rangle=0$, by definition of $\alpha_{0}^{(s)}$. Since $\alpha_{0}^{(s)}\neq 0$, then
$\left\langle\mathbf{p}_{0},\mathbf{A}^{s+1}\mathbf{p}_{1}\right\rangle$ $=\left\langle\mathbf{A}\mathbf{p}_{0},\mathbf{A}^{s}(\mathbf{r}_{1}+\beta_{0}^{(s)}\mathbf{p}_{0})\right\rangle$
$=\frac{1}{\alpha_{0}^{(s)}}\left\langle\mathbf{r}_{0}-\mathbf{r}_{1},\mathbf{A}^{s}\mathbf{r}_{1}\right\rangle+\frac{\beta_{0}^{(s)}}{\alpha_{0}^{(s)}}\left\langle\mathbf{r}_{0}-\mathbf{r}_{1},\mathbf{A}^{s}\mathbf{p}_{0}\right\rangle$
$=-\frac{1}{\alpha_{0}^{(s)}}\left\langle\mathbf{r}_{1},\mathbf{A}^{s}\mathbf{r}_{1}\right\rangle+\frac{\beta_{0}^{(s)}}{\alpha_{0}^{(s)}}\left\langle\mathbf{r}_{0},\mathbf{A}^{s}\mathbf{p}_{0}\right\rangle=0,$
by definition of $\beta_{0}^{(s)}$. Assume
$\left\langle\mathbf{r}_{i},\mathbf{A}^{s}\mathbf{r}_{j}\right\rangle$ $=0,\quad 0\leq i<j\leq k-1,$ (10a)
$\left\langle\mathbf{p}_{i},\mathbf{A}^{s+1}\mathbf{p}_{j}\right\rangle$ $=0,\quad 0\leq i<j\leq k-1,$ (10b)
for some $k$. Now we show the claim holds for $0\leq i<j\leq k$. Let $j=k$, for $i=k-1$, we have
$\left\langle\mathbf{r}_{k-1},\mathbf{A}^{s}\mathbf{r}_{k}\right\rangle$ $=\left\langle\mathbf{r}_{k-1},\mathbf{A}^{s}(\mathbf{r}_{k-1}-\alpha_{k-1}^{(s)}\mathbf{A}\mathbf{p}_{k-1})\right\rangle$
$=\left\langle\mathbf{r}_{k-1},\mathbf{A}^{s}\mathbf{r}_{k-1}\right\rangle-\alpha_{k-1}^{(s)}\left\langle(\mathbf{p}_{k-1}-\beta_{k-2}^{(s)}\mathbf{p}_{k-2})\mathbf{A}^{s+1}\mathbf{p}_{k-1}\right\rangle$
$=\left\langle\mathbf{r}_{k-1},\mathbf{A}^{s}\mathbf{r}_{k-1}\right\rangle-\alpha_{k-1}^{(s)}\left\langle\mathbf{p}_{k-1},\mathbf{A}^{s+1}\mathbf{p}_{k-1}\right\rangle=0,$
where the second last and the last equality follow from the inductive hypothesis and the definition of $\alpha_{k-1}^{(s)}$. For $i<k-1$, we have
$\left\langle\mathbf{r}_{i},\mathbf{A}^{s}\mathbf{r}_{k}\right\rangle$ $=\left\langle\mathbf{r}_{i},\mathbf{A}^{s}(\mathbf{r}_{k-1}-\alpha_{k-1}^{(s)}\mathbf{A}\mathbf{p}_{k-1})\right\rangle$
$=-\alpha_{k-1}^{(s)}\left\langle\mathbf{p}_{i}-\beta_{i-1}^{(s)}\mathbf{p}_{i-1},\mathbf{A}^{s+1}\mathbf{p}_{k-1}\right\rangle=0,$
by 10. This gives 9a. We now show 9b. For $i=k-1$, we have
$\left\langle\mathbf{p}_{k-1},\mathbf{A}^{s+1}\mathbf{p}_{k}\right\rangle$ $=\left\langle\mathbf{A}\mathbf{p}_{k-1},\mathbf{A}^{s}(\mathbf{r}_{k}+\beta_{k-1}^{(s)}\mathbf{p}_{k-1})\right\rangle$
$=\frac{1}{\alpha_{k-1}^{(s)}}\left\langle\mathbf{r}_{k-1}-\mathbf{r}_{k},\mathbf{A}^{s}\mathbf{r}_{k}\right\rangle+\beta_{k-1}^{(s)}\left\langle\mathbf{p}_{k-1},\mathbf{A}^{s+1}\mathbf{p}_{k-1}\right\rangle$
$=-\frac{1}{\alpha_{k-1}^{(s)}}\left\langle\mathbf{r}_{k},\mathbf{A}^{s}\mathbf{r}_{k}\right\rangle+\beta_{k-1}^{(s)}\left\langle\mathbf{p}_{k-1},\mathbf{A}^{s+1}\mathbf{r}_{k-1}\right\rangle$
$=-\frac{1}{\alpha_{k-1}^{(s)}}\left\langle\mathbf{r}_{k},\mathbf{A}^{s}\mathbf{r}_{k}\right\rangle+\frac{\beta_{k-1}^{(s)}}{\alpha_{k-1}^{(s)}}\left\langle\mathbf{r}_{k-1},\mathbf{A}^{s}\mathbf{r}_{k-1}\right\rangle=0,$
by 9a and the definition of $\alpha_{k-1}^{(s)}$ and $\beta_{k-1}^{(s)}$. For $i<k-1$, we have
$\left\langle\mathbf{p}_{i},\mathbf{A}^{s+1}\mathbf{p}_{k}\right\rangle$ $=\left\langle\mathbf{p}_{i},\mathbf{A}^{s+1}(\mathbf{r}_{k}+\beta_{k-1}^{(s)}\mathbf{p}_{k-1})\right\rangle$
$=\frac{1}{\alpha_{i-1}^{(s)}}\left\langle\mathbf{r}_{i}-\mathbf{r}_{i+1},\mathbf{A}^{s}\mathbf{r}_{k}\right\rangle+\beta_{k-1}^{(s)}\left\langle\mathbf{p}_{i},\mathbf{A}^{s+1}\mathbf{p}_{k-1}\right\rangle=0,$
by 9a and 10b and our induction assumptions. ∎
Lemma 11.
Let $\{\mathbf{r}_{0},\cdots,\mathbf{r}_{k-1}\}$ and $\{\mathbf{p}_{0},\cdots,\mathbf{p}_{k-1}\}$ be the set of vectors generated by Algorithm 5 up to iteration $k-1$, then, for $10\leq j<i\leq k-$ 1,
$\left\langle\mathbf{r}_{i},\mathbf{A}^{s}\mathbf{p}_{j}\right\rangle$ $=0,$ (11a)
and, for $10\leq i\leq j\leq k-$ 1,
$\left\langle\mathbf{r}_{i},\mathbf{A}^{s}\mathbf{p}_{j}\right\rangle$ $=\left\langle\mathbf{r}_{j},\mathbf{A}^{s}\mathbf{r}_{j}\right\rangle.$ (11b)
Proof.
We will prove this by induction. First, we show the base case for 11a, for $j+1=i=1$,
$\left\langle\mathbf{r}_{1},\mathbf{A}^{s}\mathbf{p}_{0}\right\rangle$ $=\left\langle\mathbf{r}_{0}-\alpha_{0}^{(s)}\mathbf{A}\mathbf{p}_{0},\mathbf{A}^{s}\mathbf{p}_{0}\right\rangle=\left\langle\mathbf{r}_{0},\mathbf{A}^{s}\mathbf{p}_{0}\right\rangle-\alpha_{0}^{(s)}\left\langle\mathbf{p}_{0},\mathbf{A}^{s+1}\mathbf{p}_{0}\right\rangle=0,$
where the last equality is by the definition of $\alpha_{0}^{(s)}$ and $\mathbf{r}_{0}$. Next, we show the base case for 11b, for $j=i=0$, we obtain the desired result trivially. Now, for $j=i=1$, we obtain $\left\langle\mathbf{r}_{1},\mathbf{A}^{s}\mathbf{p}_{1}\right\rangle=\left\langle\mathbf{r}_{1},\mathbf{A}^{s}\left(\mathbf{r}_{1}+\beta_{0}^{(s)}\mathbf{p}_{0}\right)\right\rangle=\left\langle\mathbf{r}_{1},\mathbf{A}^{s}\mathbf{r}_{1}\right\rangle$, by the previous result. For $i+1=j=1$, we have
$\left\langle\mathbf{r}_{0},\mathbf{A}^{s}\mathbf{p}_{1}\right\rangle=\left\langle\mathbf{r}_{0},\mathbf{A}^{s}\left(\mathbf{r}_{1}+\beta_{0}^{(s)}\mathbf{p}_{0}\right)\right\rangle=\left\langle\mathbf{r}_{0},\mathbf{A}^{s}\mathbf{r}_{1}\right\rangle+\beta_{0}^{(s)}\left\langle\mathbf{r}_{0},\mathbf{A}^{s}\mathbf{p}_{0}\right\rangle=\left\langle\mathbf{r}_{1},\mathbf{A}^{s}\mathbf{r}_{1}\right\rangle,$
where the last line is by the definition of $\beta_{0}^{(s)}$ and 9a. Now, assume
$\left\langle\mathbf{r}_{i},\mathbf{A}^{s}\mathbf{p}_{j}\right\rangle$ $=0,\quad 0\leq j<i\leq k-1,$
$\left\langle\mathbf{r}_{i},\mathbf{A}^{s}\mathbf{p}_{j}\right\rangle$ $=\left\langle\mathbf{r}_{j},\mathbf{A}^{s}\mathbf{r}_{j}\right\rangle,\quad 0\leq i\leq j\leq k-1$
for some $k$. We now show the claim for $k+1$. Let $i=k$, for $j=k-1$, we have
$\left\langle\mathbf{r}_{k},\mathbf{A}^{s}\mathbf{p}_{k-1}\right\rangle$ $=\left\langle\mathbf{r}_{k-1}-\alpha_{k-1}^{(s)}\mathbf{Ap}_{k-1},\mathbf{A}^{s}\mathbf{p}_{k-1}\right\rangle$
$=\left\langle\mathbf{r}_{k-1},\mathbf{A}^{s}\mathbf{p}_{k-1}\right\rangle-\alpha_{k-1}^{(s)}\left\langle\mathbf{p}_{k-1},\mathbf{A}^{s+1}\mathbf{p}_{k-1}\right\rangle=0,$
by the definition of $\alpha_{k-1}^{(s)}$ and the inductive hypothesis. For $j<k-1$, we get
$\left\langle\mathbf{r}_{k},\mathbf{A}^{s}\mathbf{p}_{j}\right\rangle$ $=\left\langle\mathbf{r}_{k-1}-\alpha_{k-1}^{(s)}\mathbf{Ap}_{k-1},\mathbf{A}^{s}\mathbf{p}_{j}\right\rangle$
$=\left\langle\mathbf{r}_{k-1},\mathbf{A}^{s}\mathbf{p}_{j}\right\rangle-\alpha_{k-1}^{(s)}\left\langle\mathbf{p}_{k-1},\mathbf{A}^{s+1}\mathbf{p}_{j}\right\rangle=0,$
where the first term and the second term each evaluate to zero by the inductive hypothesis and 9b, respectively. This gives us 11a. Now, let $j=k$, for $i=k$, we have
$\left\langle\mathbf{r}_{k},\mathbf{A}^{s}\mathbf{p}_{k}\right\rangle=\left\langle\mathbf{r}_{k},\mathbf{A}^{s}\left(\mathbf{r}_{k}+\beta_{k-1}^{(s)}\mathbf{p}_{k-1}\right)\right\rangle=\left\langle\mathbf{r}_{k},\mathbf{A}^{s}\mathbf{r}_{k}\right\rangle,$
by 11a. For $i<k$, we have
$\left\langle\mathbf{r}_{i},\mathbf{A}^{s}\mathbf{p}_{k}\right\rangle$ $=\left\langle\mathbf{r}_{i},\mathbf{A}^{s}\left(\mathbf{r}_{k}+\beta_{k-1}^{(s)}\mathbf{p}_{k-1}\right)\right\rangle=\left\langle\mathbf{r}_{i},\mathbf{A}^{s}\mathbf{r}_{k}\right\rangle+\beta_{k-1}^{(s)}\left\langle\mathbf{r}_{i},\mathbf{A}^{s}\mathbf{p}_{k-1}\right\rangle$
$=\beta_{k-1}^{(s)}\left\langle\mathbf{r}_{k-1},\mathbf{A}^{s}\mathbf{r}_{k-1}\right\rangle=\left\langle\mathbf{r}_{k},\mathbf{A}^{s}\mathbf{r}_{k}\right\rangle,$
where the last line is by the definition of $\beta_{0}^{(s)}$, and the second equality due to 9a and the inductive hypothesis, respectively. This gives us 11b. ∎
Appendix B Further Results
In this section, we gather a few results that are used in the proofs throughout the paper. We first provide a simple yet useful result that holds in general regardless of consistency.
Lemma 12.
Under Assumption 1, in Algorithm 5 we have $(\mathbf{I}-\mathbf{A}^{\dagger}\mathbf{A})\mathbf{p}_{k-1}=c^{(s)}_{k-1}\mathbf{b}^{\perp}$ for $1\leq k\leq g$, where $c^{(s)}_{k-1}=1+\beta^{(s)}_{k-2}c^{(s)}_{k-2},\;k\geq 1$, with $c^{(s)}_{0}=1$.
Proof.
We first note that, for any $\mathbf{x}$, $(\mathbf{I}-\mathbf{A}^{\dagger}\mathbf{A})\mathbf{r}=(\mathbf{I}-\mathbf{A}^{\dagger}\mathbf{A})(\mathbf{b}-\mathbf{A}\mathbf{x})=\mathbf{b}^{\perp}$ where we have used the fact that, for the symmetric matrix $\mathbf{A}$, we have $\mathbf{A}^{\dagger}\mathbf{A}=\mathbf{A}\mathbf{A}^{\dagger}$. The proof follows an inductive argument. The result trivially holds for $k=1$. Suppose the claim holds for any $1\leq k<g$. For $k+1$, we have
$(\mathbf{I}-\mathbf{A}^{\dagger}\mathbf{A})\mathbf{p}_{k}$ $=(\mathbf{I}-\mathbf{A}^{\dagger}\mathbf{A})(\mathbf{r}_{k}+\beta_{k-1}^{(s)}\mathbf{p}_{k-1})=\mathbf{b}^{\perp}+\beta^{(s)}_{k-1}(\mathbf{I}-\mathbf{A}^{\dagger}\mathbf{A})\mathbf{p}_{k-1}$
$=\left(1+\beta^{(s)}_{k-1}c^{(s)}_{k-1}\right)\mathbf{b}^{\perp}.$
The next result, which is crucial for several proofs, particularly for the derivations in Appendix C, states that the contribution of the null space of $\mathbf{A}$ to $\mathbf{b}$, namely $\mathbf{b}^{\perp}$, is included in the Krylov subspace only once the grade of $\mathbf{b}$ with respect to $\mathbf{A}$ is reached.
Lemma 13.
Let $\mathbf{b}\notin\text{Range}(\mathbf{A})$. Then, $\mathbf{b}^{\perp}\in\mathcal{K}_{k}(\mathbf{A},\mathbf{b})$ if and only if $k\geq g$.
Proof.
First, we note that since $\mathbf{b}\notin\text{Range}(\mathbf{A})$, we must have $\mathbf{b}\neq\mathbf{0}$ and $\mathbf{b}^{\perp}\neq\mathbf{0}$. Suppose $\mathbf{b}^{\perp}\in\mathcal{K}_{k}(\mathbf{A},\mathbf{b})$ but $k<g$. Recall that in this case $\{\mathbf{b},\mathbf{A}\mathbf{b},\cdots,\mathbf{A}^{k-1}\mathbf{b},\mathbf{A}^{k}\mathbf{b}\}$ are linearly independent. We can write
$\mathbf{b}^{\perp}$ $=a_{0}\mathbf{b}+a_{1}\mathbf{Ab}+\cdots+a_{k-1}\mathbf{A}^{k-1}\mathbf{b},$
which gives
$\mathbf{0}=\mathbf{Ab}^{\perp}$ $=a_{0}\mathbf{Ab}+a_{1}\mathbf{A}^{2}\mathbf{b}+\cdots+a_{k-1}\mathbf{A}^{k}\mathbf{b}\in\mathcal{K}_{k+1}(\mathbf{A},\mathbf{b}).$
Since not all the coefficients $a_{i}$, $10\leq i\leq k-$ 1, are zero, the vectors $\mathbf{Ab},\mathbf{A}^{2}\mathbf{b},\cdots,\mathbf{A}^{k}\mathbf{b}$ must be linearly dependent, which leads to a contradiction. Conversely, suppose $k\geq g$. We can write $\mathbf{0}=a_{0}\mathbf{b}+a_{1}\mathbf{Ab}+\cdots+a_{g}\mathbf{A}^{g}\mathbf{b}$ where not all coefficients $a_{i},0\leq i\leq g$, are zero. By assumption $\mathbf{b}\notin\text{Range}(\mathbf{A})$, which implies that $a_{0}=0$. This gives
$\mathbf{0}$ $=a_{1}\mathbf{Ab}+a_{2}\mathbf{A}^{2}\mathbf{b}+\cdots+a_{g}\mathbf{A}^{g}\mathbf{b}=\mathbf{A}(a_{1}\mathbf{b}+a_{2}\mathbf{A}\mathbf{b}+\cdots+a_{g}\mathbf{A}^{g-1}\mathbf{b})=\mathbf{Ac},$
where $\mathbf{c}=a_{1}\mathbf{b}+a_{2}\mathbf{A}\mathbf{b}+\cdots+a_{g}\mathbf{A}^{g-1}\mathbf{b}\in\mathcal{K}_{g}(\mathbf{A},\mathbf{b})$. Hence, $\mathbf{0}\neq\mathbf{c}\in\text{Null}(\mathbf{A})$. Clearly, $a_{1}\neq 0$, as otherwise $\mathbf{c}\in\text{Range}(\mathbf{A})$, which implies $\mathbf{c}=\mathbf{0}$. Now, from $\mathbf{b}=\tilde{\mathbf{b}}+\mathbf{b}^{\perp}$, we get
$\frac{1}{a_{1}}\mathbf{c}-\mathbf{b}^{\perp}=\tilde{\mathbf{b}}+\frac{1}{a_{1}}\left(a_{2}\mathbf{A}\mathbf{b}+\cdots+a_{g}\mathbf{A}^{g-1}\mathbf{b}\right).$
The left-hand side is a vector in the null space of $\mathbf{A}$, where as the right-hand side is in the range of $\mathbf{A}$. This equality can only hold if both sides are zero vectors. Consequently, we have $\mathbf{b}^{\perp}=\mathbf{c}/a_{1}\in\mathcal{K}_{g}(\mathbf{A},\mathbf{b})$
The next result states that if the subspace $\mathcal{K}_{k}(\mathbf{A},\mathbf{b})$ contains any normal solution, it must be of a specific format.
Lemma 14.
If $\mathbf{A}^{\dagger}\mathbf{b}+\mathbf{y}\in\mathcal{K}_{k}(\mathbf{A},\mathbf{b})$ for some $\mathbf{y}\in\text{Null}(\mathbf{A})$, then $\mathbf{y}=a(\mathbf{I}-\mathbf{A}\mathbf{A}^{\dagger})\mathbf{b}$ for some $a\in$.
Proof.
Suppose $\mathbf{A}^{\dagger}\mathbf{b}+\mathbf{y}\in\mathcal{K}_{k}(\mathbf{A},\mathbf{b})$ for some $\mathbf{y}\in\text{Null}(\mathbf{A})$. We have $\mathbf{A}^{\dagger}\mathbf{b}+\mathbf{y}=a_{0}\mathbf{b}+a_{1}\mathbf{Ab}+\cdots+a_{k-1}\mathbf{A}^{k-1}\mathbf{b}$, which implies $\mathbf{y}=(\mathbf{I}-\mathbf{A}\mathbf{A}^{\dagger})\mathbf{y}=a_{0}(\mathbf{I}-\mathbf{A}\mathbf{A}^{\dagger})\mathbf{b}$. ∎
The next result indicates that the underlying Krylov subspace can contain a normal solution only if it is large enough. This result is likely already known; however, we could not pinpoint a source to reference it. Therefore, we include it here for completeness.
Lemma 15.
  • Let $\mathbf{b}\notin\text{Range}(\mathbf{A})$. Then $\mathbf{A}^{\dagger}\mathbf{b}+a(\mathbf{I}-\mathbf{A}\mathbf{A}^{\dagger})\mathbf{b}\in\mathcal{K}_{k}(\mathbf{A},\mathbf{b})$ for some $a\in$ if and only if $k\geq g-1$.
  • Let $\mathbf{b}\in\text{Range}(\mathbf{A})$. Then $\mathbf{A}^{\dagger}\mathbf{b}\in\mathcal{K}_{k}(\mathbf{A},\mathbf{b})$ if and only if $k\geq g$.
Proof.
  • Suppose $\mathbf{b}\notin\text{Range}(\mathbf{A})$, which in particular implies $\mathbf{b}\neq\mathbf{0}$ and $\mathbf{b}^{\perp}\neq\mathbf{0}$. Let $k\geq g-1$. By Lemma 13, we know that $\mathbf{b}^{\perp}\in\mathcal{K}_{g}(\mathbf{A},\mathbf{b})$. We write, $-\mathbf{b}^{\perp}=a_{0}\mathbf{b}+a_{1}\mathbf{Ab}+\cdots+a_{g-1}\mathbf{A}^{g-1}\mathbf{b}$, which can only hold if $a_{0}=-1$. Hence,
    $\tilde{\mathbf{b}}$ $=a_{1}\mathbf{Ab}+a_{2}\mathbf{A}^{2}\mathbf{b}+\cdots+a_{g-1}\mathbf{A}^{g-1}\mathbf{b}=\mathbf{A}(a_{1}\mathbf{b}+a_{2}\mathbf{A}\mathbf{b}+\cdots+a_{g-1}\mathbf{A}^{g-2}\mathbf{b})$
    $=\mathbf{A}(\underbrace{a_{1}\tilde{\mathbf{b}}+\cdots+a_{g-1}\mathbf{A}^{g-2}\mathbf{b}}_{\triangleq\,\,\mathbf{c}}+\underbrace{a_{1}\mathbf{b}^{\perp}}_{\triangleq\,\,\mathbf{c}^{\prime}}).$
    From $\mathbf{c}\in\text{Range}(\mathbf{A})$, it follows that
    $\mathbf{A}^{\dagger}\mathbf{b}=\mathbf{A}^{\dagger}\tilde{\mathbf{b}}=\mathbf{A}^{\dagger}\mathbf{A}\mathbf{c}=\mathbf{c}.$
    Now, since $\mathbf{c}+\mathbf{c}^{\prime}\in\mathcal{K}_{g-1}(\mathbf{A},\mathbf{b})\subseteq\mathcal{K}_{g}(\mathbf{A},\mathbf{b})$, we get the desired result. Conversely, assume that for some $k<g-1$, $\mathcal{K}_{k}(\mathbf{A},\mathbf{b})$ contains a normal solution, i.e., $\mathbf{A}^{\dagger}\mathbf{b}+(\mathbf{I}-\mathbf{AA}^{\dagger})\mathbf{y}\in\mathcal{K}_{k}(\mathbf{A},\mathbf{b})$ for some $\mathbf{y}$ and $k<g-1$. We can write
    $\mathbf{A}^{\dagger}\mathbf{b}+(\mathbf{I}-\mathbf{AA}^{\dagger})\mathbf{y}$ $=a_{0}\mathbf{b}+a_{1}\mathbf{Ab}+\cdots+a_{k-1}\mathbf{A}^{k-1}\mathbf{b},$
    which gives
    $\tilde{\mathbf{b}}=\mathbf{A}\mathbf{A}^{\dagger}\mathbf{b}=\mathbf{A}(\mathbf{A}^{\dagger}\mathbf{b}+(\mathbf{I}-\mathbf{AA}^{\dagger})\mathbf{y})$ $=a_{0}\mathbf{Ab}+a_{1}\mathbf{A}^{2}\mathbf{b}+\cdots+a_{k-1}\mathbf{A}^{k}\mathbf{b}.$
    Hence,
    $-\mathbf{b}^{\perp}$ $=-\mathbf{b}+a_{0}\mathbf{Ab}+a_{1}\mathbf{A}^{2}\mathbf{b}+\cdots+a_{k-1}\mathbf{A}^{k}\mathbf{b}\in\mathcal{K}_{k+1}(\mathbf{A},\mathbf{b}),$
    which contradicts Lemma 13. So, $k\geq g-1$.
  • Suppose $\mathbf{b}\in\text{Range}(\mathbf{A})$. If $\mathbf{b}=\mathbf{0}$, then the results is trivial. Assume $\mathbf{b}\neq\mathbf{0}$ and $\mathbf{A}^{\dagger}\mathbf{b}\in\mathcal{K}_{k}(\mathbf{A},\mathbf{b})$ but $k<g$. We can write
    $\mathbf{A}^{\dagger}\mathbf{b}$ $=a_{0}\mathbf{b}+a_{1}\mathbf{Ab}+\cdots+a_{k-1}\mathbf{A}^{k-1}\mathbf{b},$
    which implies
    $\mathbf{b}=\mathbf{AA}^{\dagger}\mathbf{b}$ $=a_{0}\mathbf{Ab}+a_{1}\mathbf{A}^{2}\mathbf{b}+\cdots+a_{k-1}\mathbf{A}^{k}\mathbf{b}.$
    However, this leads to a contradiction, since $\mathbf{b},\mathbf{Ab},\cdots,\mathbf{A}^{k}\mathbf{b}$ are linearly independent vectors for $k\leq g-1$. So, we must have $k\geq g$. Conversely, let $k\geq g$ and consider $\mathcal{K}_{g+1}(\mathbf{A},\mathbf{b})$. Since the vectors $\mathbf{b},\mathbf{Ab},\cdots,\mathbf{A}^{g}\mathbf{b}$ are linearly dependent, there exist coefficients $a_{0},a_{1},\ldots,a_{g}$, not all zero, such that $\mathbf{0}=-a_{0}\mathbf{b}+a_{1}\mathbf{Ab}+\cdots+a_{g}\mathbf{A}^{g}\mathbf{b}$. In particular, we must have $a_{0}\not=0$. Otherwise, applying $\mathbf{A}^{\dagger}$ twice to both sides yields
    $-a_{1}\mathbf{A}^{\dagger}\mathbf{b}=a_{2}\mathbf{b}+a_{3}\mathbf{Ab}+\cdots+a_{g}\mathbf{A}^{g-2}\mathbf{b}\in\mathcal{K}_{g-1}(\mathbf{A},\mathbf{b}).$
    However, by the first part of the present proof, this cannot occur. Hence, $a_{0}\neq 0$, and we can write
    $a_{0}\mathbf{A}^{\dagger}\mathbf{b}=a_{1}\mathbf{b}+\cdots+a_{g}\mathbf{A}^{g-1}\mathbf{b},$
    which implies $\mathbf{A}^{\dagger}\mathbf{b}\in\mathcal{K}_{g}(\mathbf{A},\mathbf{b})$.
Lemma 16.
Let $\mathbf{b}\in\text{Range}(\mathbf{A})$. Under Assumption 1, in Algorithm 5, we have $\mathbf{r}_{g}=\mathbf{0}$.
Proof.
We first note that when $\mathbf{b}\in\text{Range}(\mathbf{A})$, for the pseudo-inverse solution $\mathbf{A}^{\dagger}\mathbf{b}$, we have $\mathbf{r}^{\star}=\mathbf{b}-\mathbf{A}\mathbf{A}^{\dagger}\mathbf{b}=\mathbf{0}$. Let $\mathbf{x}^{\star}\in\mathcal{K}_{g}(\mathbf{A},\mathbf{b})$ be the pseudo-inverse solution from Lemma 15. Using Lemma 9, we can express $\mathbf{x}^{\star}$ as
$\mathbf{x}^{\star}$ $=\sum_{i=1}^{g}\alpha_{i-1}^{(s)}\mathbf{p}_{i-1},$
for some coefficients $\alpha_{i-1}^{(s)}$’s. It suffices to show that $\alpha_{i}^{(s)}$’s correspond to those from Algorithm 5. From $\mathbf{0}=\mathbf{r}^{\star}=\mathbf{b}-\mathbf{A}\mathbf{x}^{\star}=\mathbf{b}-\sum_{i=1}^{g}\alpha_{i-1}^{(s)}\mathbf{Ap}_{i-1}$, we have
$\mathbf{0}=\mathbf{A}^{s}\mathbf{r}^{\star}$ $=\mathbf{A}^{s}\mathbf{b}-\sum_{i=1}^{g}\alpha_{i-1}^{(s)}\mathbf{A}^{s+1}\mathbf{p}_{i-1}.$
Multiplying $\mathbf{p}_{j}$, for each $10\leq j\leq g-$ 1, from the left, using Lemma 10 as well as $\mathbf{p}_{j}=\mathbf{r}_{j}+\beta_{j-1}\mathbf{p}_{j-1}$, gives $\alpha_{i-1}^{(s)}=\left\langle\mathbf{r}_{i-1},\mathbf{A}^{s}\mathbf{r}_{i-1}\right\rangle/\left\langle\mathbf{p}_{i-1},\mathbf{A}^{s+1}\mathbf{p}_{i-1}\right\rangle$, which are precisely the step-sizes calculated within Algorithm 5. Thus, $\mathbf{r}_{g}=\mathbf{r}^{\star}=\mathbf{0}$. ∎
Appendix C Orthogonal-Residual and Inconsistent Symmetric Systems
In this section, we provide further insights into why CG fails to converge for inconsistent systems. Traditionally, CG is viewed as a specific algorithmic implementation of the Orthogonal-Residual (OR) framework [^8] Also referred to as the Full Orthogonalization Method (FOM) in other works [40]. when applied to PD matrices. More generally, OR enforces the following projection conditions:
$\mathbf{x}_{k}$ $\in\mathcal{K}_{k}(\mathbf{A},\mathbf{b}),$ (12a)
$\mathbf{r}_{k}$ $\perp\mathcal{K}_{k}(\mathbf{A},\mathbf{b}),$ (12b)
where $\mathbf{A}$ is a square but not necessarily symmetric matrix. In the special case where $\mathbf{A}$ is PD, the Arnoldi process to construct a basis for $\mathcal{K}_{k}(\mathbf{A},\mathbf{b})$ reduces to the Lanczos process, and OR can be algorithmically implemented as Algorithm 1, i.e., the CG algorithm (see [40] for further details).
Remark 9.
Under this interpretation, we distinguish between the framework and its algorithmic implementations. For example, the Minimum-Residual (MR) framework is defined as a projection framework that enforces $\mathbf{x}_{k}\in\mathcal{K}_{k}(\mathbf{A},\mathbf{b})$ and $\mathbf{r}_{k}\perp\mathbf{A}\mathcal{K}_{k}(\mathbf{A},\mathbf{b})$. One particular algorithmic implementation of the MR framework is GMRES. For symmetric matrices, other implementations include the CR and MINRES algorithms; see Table 3 for an overview.
Krylov Subspace Projection Framework Algorithmic implementations of the framework
$\mathbf{A}$ is square $\mathbf{A}$ is PD $\mathbf{A}$ is PSD $\mathbf{A}$ is indefinite
$\mathbf{b}\in\text{R}(\mathbf{A})$ $\mathbf{b}\notin\text{R}(\mathbf{A})$ $\mathbf{b}\in\text{R}(\mathbf{A})$ $\mathbf{b}\notin\text{R}(\mathbf{A})$
MR framework GMRES CR, CR, CR, CR*, CR*,
$\mathbf{r}_{k}\perp\mathbf{A}\mathcal{K}_{k}(\mathbf{A},\mathbf{b})$ MINRES MINRES MINRES MINRES MINRES
OR framework FOM[40] CG, CG, CG*,
$\mathbf{r}_{k}\perp\mathcal{K}_{k}(\mathbf{A},\mathbf{b})$ LanczosCG[4] LanczosCG LanczosCG*
Table 3: Range$(\mathbf{A})$ is abbreviated as R$(\mathbf{A})$ in this table. An overview of the MR and OR frameworks, along with their respective algorithmic implementations under various conditions. The symbols $\ast$ and ✗  indicate, respectively, the necessity of Assumption 1 and the nonexistence of such algorithms (see Theorem 6).
As demonstrated in Section 3.1, CG fails to converge to a normal solution for inconsistent systems. Given that CG is derived from the OR framework, one might ask whether an alternative algorithmic implementation of OR for symmetric matrices could converge to a normal solution under inconsistent systems. We establish that no such implementation exists. To demonstrate this, we restate Lemma 17, originally presented in [4, Proposition 2.7], and provide a proof here for completeness.
Lemma 17.
Let $\mathbf{b}\in^{d}$ and $\mathbf{A}\in^{d\times d}$ be symmetric matrix. Consider the Lanczos process, which generates the orthonormal matrix $\mathbf{V}_{g}\in\mathbb{R}^{d\times g}$ and the tridiagonal matrix $\mathbf{T}_{g}\in\mathbb{R}^{g\times g}$, such that $\mathbf{AV}_{g}=\mathbf{V}_{g}\mathbf{T}_{g}$. We have $\mathbf{b}\notin\text{Range}(\mathbf{A})$ if and only if $\mathbf{T}_{g}$ is singular.
Proof.
First, we note that since $\mathbf{b}\notin\text{Range}(\mathbf{A})$, we must have $\mathbf{b}\neq\mathbf{0}$ and $\mathbf{b}^{\perp}\neq\mathbf{0}$. By the Lanczos process and Lemma 13, given that $\mathbf{b}^{\perp}\in\mathcal{K}_{g}(\mathbf{A},\mathbf{b})=\text{Span}(\mathbf{V}_{g})$, there exists a nonzero vector $\mathbf{c}\in\mathbb{R}^{g}$ such that $\mathbf{V}_{g}\mathbf{c}=\mathbf{b}^{\perp}$. Furthermore, since
$\mathbf{T}_{g}\mathbf{c}=\mathbf{V}_{g}^{\top}\mathbf{AV}_{g}\mathbf{c}=\mathbf{V}_{g}^{\top}\mathbf{A}\mathbf{b}^{\perp}=\mathbf{0},$
it follows that $\mathbf{T}_{g}$ is singular. Conversely, let $\mathbf{c}\neq\mathbf{0}$ be a vector in $\text{Null}(\mathbf{T}_{g})$. Then, $\mathbf{AV}_{g}\mathbf{c}=\mathbf{V}_{g}\mathbf{T}_{g}\mathbf{c}=\mathbf{0}$. Hence, we must have $\mathbf{0}\neq\mathbf{V}_{g}\mathbf{c}\in\text{Null}(\mathbf{A})$. As a result, given that $\text{Span}(\mathbf{V}_{g})=\text{Span}([\mathbf{b},\mathbf{Ab},\dots,\mathbf{A}^{g-1}\mathbf{b}])$, we conclude that $\mathbf{V}_{g}\mathbf{c}=\alpha\mathbf{b}^{\perp}$ for some nonzero constant $\alpha$. ∎
We are now ready to state and prove the main result of this section.
Theorem 6.
Any algorithm derived from the Orthogonal-Residual (OR) framework for symmetric matrices cannot converge to a solution of the normal equation under inconsistent systems.
Proof.
By Lemma 15, a normal solution can only exist in either $\mathcal{K}_{g-1}(\mathbf{A},\mathbf{b})$ or $\mathcal{K}_{g}(\mathbf{A},\mathbf{b})$. The proof involves demonstrating that any algorithm based on the OR framework cannot find a normal solution in either of these two subspaces. Assume the contrary that , i.e., there exists an algorithm based on the OR framework 12 generating $\mathbf{x}_{k}=\mathbf{A}^{\dagger}\mathbf{b}+a(\mathbf{I}-\mathbf{A}\mathbf{A}^{\dagger})\mathbf{b}\in\mathcal{K}_{k}(\mathbf{A},\mathbf{b})$ for some $a$, where $k=g-1$ or $k=g$ (recall that by Lemma 14, this can be the only form of a normal solution in these spaces).
  1. (i)
    First, consider the former case of $k=g-1$. Let $\mathbf{V}_{k}\in\mathbb{R}^{d\times k}$ be the orthogonal matrix whose columns are the basis vectors for $\mathcal{K}_{k}(\mathbf{A},\mathbf{b})$ such that $\mathbf{AV}_{k}=\mathbf{V}_{k}\mathbf{T}_{k}+\beta_{k}\mathbf{v}_{k+1}\mathbf{e}_{k}^{\top}$ where $\mathbf{T}_{k}\in\mathbb{R}^{k\times k}$ is a symmetric tridiagonal matrix. Such construction is guaranteed by the Lanczos process. From 12a, we can write $\mathbf{x}_{g-1}=\mathbf{V}_{g-1}\mathbf{y}$, which implies,
    $\mathbf{0}=\mathbf{V}_{g-1}^{\top}\mathbf{r}_{g-1}$ $=\mathbf{V}_{g-1}^{\top}(\mathbf{b}-\mathbf{A}\mathbf{V}_{g-1}\mathbf{y})=\|\mathbf{b}\|\mathbf{e}_{1}-\mathbf{T}_{g-1}\mathbf{y},$
    where the first equality follows from 12b. Our proof now consists of showing that no vector $\mathbf{y}$ can satisfy the above equality.
    • Suppose $\mathbf{T}_{g-1}$ is nonsingular. We have $\mathbf{y}=\mathbf{T}_{g-1}^{-1}\|\mathbf{b}\|\mathbf{e}_{1}$. Since $\mathbf{x}_{g-1}$ is a normal solution, i.e., $\mathbf{A}\mathbf{r}_{g-1}=\mathbf{0}$, it follows that
      $\mathbf{0}$ $=\mathbf{A}(\mathbf{b}-\mathbf{Ax}_{g-1})=\mathbf{A}(\mathbf{b}-\mathbf{AV}_{g-1}\mathbf{y})=\mathbf{A}\left(\mathbf{b}-\begin{bmatrix}\mathbf{V}_{g-1}&\mathbf{v}_{g}\end{bmatrix}\begin{bmatrix}\mathbf{T}_{g-1}\\ \beta_{g}\mathbf{e}_{g-1}^{\top}\end{bmatrix}\mathbf{y}\right)$
      $=\mathbf{A}(\mathbf{b}-(\mathbf{V}_{g-1}\mathbf{T}_{g-1}+\beta_{g}\mathbf{v}_{g}\mathbf{e}_{g-1}^{\top})\mathbf{y})=\mathbf{A}(\mathbf{b}-(\mathbf{V}_{g-1}\mathbf{T}_{g-1}+\beta_{g}\mathbf{v}_{g}\mathbf{e}_{g-1}^{\top})\mathbf{T}_{g-1}^{-1}\|\mathbf{b}\|\mathbf{e}_{1})$
      $=\mathbf{A}(\mathbf{b}-(\mathbf{b}+\beta_{g}\mathbf{v}_{g}\mathbf{e}_{g-1}^{\top}\mathbf{T}_{g-1}^{-1}\|\mathbf{b}\|\mathbf{e}_{1}))=-\beta_{g}\mathbf{A}\mathbf{v}_{g}\mathbf{e}_{g-1}^{\top}\mathbf{T}_{g-1}^{-1}\|\mathbf{b}\|\mathbf{e}_{1}.$
      Recall the construction of the vector $\mathbf{v}_{g}$ from the three-term recurrence of the Lanczos process,
      $\mathbf{Av}_{g-1}-\beta_{g-1}\mathbf{v}_{g-2}-\alpha_{g-1}\mathbf{v}_{g-1}=\beta_{g}\mathbf{v}_{g}.$
      Since all the vectors on the left-hand side of the above equality are non-zero and linearly independent, we have $\mathbf{v}_{g}^{\top}\mathbf{Av}_{g-1}=\beta_{g}\not=0$ by the orthogonality of the vectors $\mathbf{v}_{i}$. Together, we obtain,
      00 0 $=\mathbf{v}_{g-1}^{\top}\mathbf{0}=-\beta_{g}\mathbf{v}_{g-1}^{\top}\mathbf{A}\mathbf{v}_{g}\mathbf{e}_{g-1}^{\top}\mathbf{T}_{g-1}^{-1}\|\mathbf{b}\|\mathbf{e}_{1}=-\beta_{g}^{2}\|\mathbf{b}\|\mathbf{e}_{g-1}^{\top}\mathbf{T}_{g-1}^{-1}\mathbf{e}_{1}$
      $=-\beta_{g}^{2}\|\mathbf{b}\|\mathbf{e}_{g-1}^{\top}\mathbf{c}=-\beta_{g}^{2}\|\mathbf{b}\|c_{g-1},$
      where $\mathbf{c}\triangleq\mathbf{T}_{g-1}^{-1}\mathbf{e}_{1}$. Since $\beta_{g}^{2}\|\mathbf{b}\|\neq 0$, we must have $c_{g-1}=0$. In this case, we have
      $\mathbf{e}_{1}=\mathbf{T}_{g-1}\mathbf{c}$ $=\begin{bmatrix}\alpha_{1}&\beta_{2}&&&\\ \beta_{2}&\alpha_{2}&\beta_{3}&&\\ &&&\ddots&\beta_{g-1}\\ &&&\beta_{g-1}&\alpha_{g-1}\end{bmatrix}\begin{bmatrix}c_{1}\\ c_{2}\\ \vdots\\ c_{g-1}\end{bmatrix}$
      $=\begin{bmatrix}c_{1}\alpha_{1}+c_{2}\beta_{2}\\ \vdots\\ c_{g-3}\beta_{g-2}+c_{g-2}\alpha_{g-2}+c_{g-1}\beta_{g-1}\\ c_{g-2}\beta_{g-1}+c_{g-1}\alpha_{g-1}\end{bmatrix}=\begin{bmatrix}c_{1}\alpha_{1}+c_{2}\beta_{2}\\ \vdots\\ c_{g-3}\beta_{g-2}+c_{g-2}\alpha_{g-2}\\ c_{g-2}\beta_{g-1}\end{bmatrix}.$
      By the construction of the Lanczos process, the constants $\beta_{2},\cdots,\beta_{g-1}$ are non-zero. This implies that $c_{g-2}=c_{g-3}=\cdots=c_{1}=0$, and hence $\mathbf{c}=\mathbf{0}$. However, since $\mathbf{T}_{g-1}$ is assumed to be nonsingular, $\mathbf{T}_{g-1}^{-1}\mathbf{e}_{1}$ cannot be a zero vector, which leads to a contradiction.
    • So now suppose $\mathbf{T}_{g-1}$ is singular, which in particular implies that it has a nontrivial null-space. We need to show that $\mathbf{e}_{1}\notin\text{Range}(\mathbf{T}_{g-1})$. Suppose to the contrary that $\mathbf{e}_{1}\in\text{Range}(\mathbf{T}_{g-1})$, or equivalently, for all non-zero $\mathbf{c}\in\text{Null}(\mathbf{T}_{g-1})$, we have $\langle\mathbf{e}_{1},\mathbf{c}\rangle=c_{1}=0$. Take any non-zero $\mathbf{c}\in\text{Null}(\mathbf{T}_{g-1})$. It follows that
      $\mathbf{0}=\mathbf{T}_{g-1}\mathbf{c}=\begin{bmatrix}\alpha_{1}&\beta_{2}&&&\\ \beta_{2}&\alpha_{2}&\beta_{3}&&\\ &&&\ddots&\beta_{g-1}\\ &&&\beta_{g-1}&\alpha_{g-1}\end{bmatrix}\begin{bmatrix}c_{1}\\ c_{2}\\ \vdots\\ c_{g-1}\end{bmatrix}=\begin{bmatrix}c_{1}\alpha_{1}+c_{2}\beta_{2}\\ c_{1}\beta_{2}+c_{2}\alpha_{2}+c_{3}\beta_{3}\\ \vdots\\ c_{g-2}\beta_{g-1}+c_{g-1}\alpha_{g-1}\end{bmatrix}.$
      Again, by the constriction of the Lanczos process, $\beta_{2},\cdots,\beta_{g-1}$ are non-zero constants, and hence if $\langle\mathbf{e}_{1},\mathbf{c}\rangle=c_{1}=0$, then $c_{2}=c_{3}=\cdots=c_{g-1}=0$. This implies $\mathbf{c}=\mathbf{0}$, contradicting the non-zero assumption on $\mathbf{c}$. Hence, it follows that there must exists a non-zero $\mathbf{c}\in\text{Null}(\mathbf{T}_{g-1})$ such that $\langle\mathbf{e}_{1},\mathbf{c}\rangle\neq 0$, i.e., $\mathbf{e}_{1}\notin\text{Range}(\mathbf{T}_{g-1})$.
    The above argument shows that no method based on the OR framework for symmetric matrices under inconsistent systems can find a normal solution in the subspace $\mathcal{K}_{g-1}(\mathbf{A},\mathbf{b})$.
  2. (ii)
    Now, we consider the latter case of $k=g$. By 12a, $\mathbf{x}_{g}=\mathbf{V}_{g}\mathbf{y}$, and by 12b,
    $\mathbf{0}=\mathbf{V}_{g}^{\top}\mathbf{r}_{g}$ $=\mathbf{V}_{g}^{\top}(\mathbf{b}-\mathbf{AV}_{g}\mathbf{y})=\|\mathbf{b}\|\mathbf{e}_{1}-\mathbf{V}_{g}^{\top}\mathbf{AV}_{g}\mathbf{y}=\|\mathbf{b}\|\mathbf{e}_{1}-\mathbf{T}_{g}\mathbf{y}.$
    However, by Lemma 17, $\mathbf{T}_{g}$ is singular. So, a similar argument as above can establish that no vector $\mathbf{y}$ can satisfy the above equality, i.e., the conditions 12a and 12b cannot be simultaneously satisfied in the subspace $\mathcal{K}_{g}(\mathbf{A},\mathbf{b})$.